### v1

# Randomization Tests for Conditional Group Symmetry

Abstract
Symmetry plays a central role in the sciences, machine learning, and statistics. While statistical tests for the presence of distributional invariance with respect to groups have a long history, tests for conditional symmetry in the form of equivariance or conditional invariance are absent from the literature. This work initiates the study of nonparametric randomization tests for symmetry (invariance or equivariance) of a conditional distribution under the action of a specified locally compact group. We develop a general framework for randomization tests with finite-sample Type I error control and, using kernel methods, implement tests with finite-sample power lower bounds. We also describe and implement approximate versions of the tests, which are asymptotically consistent. We study their properties empirically on synthetic examples, and on applications to testing for symmetry in two problems from high-energy particle physics.
62G10,
62H15,
62H05,
62P35,
equivariance,
invariance,
symmetry,
hypothesis test,
randomization test,
conditional independence,
keywords:
[class=MSC]
keywords:
[…]
, and
1 Introduction
Symmetry plays a central role in science and engineering, and appears in numerous statistical problems both classical (Lehmann and Casella, 1998; Lehmann and Romano, 2005) and modern (e.g., Cohen, Geiger and Weiler, 2019; Chen, Dobriban and Lee, 2020; Huang, Orbanz and Austern, 2022; Bronstein et al., 2021). There is an extensive literature on hypothesis tests for symmetry in the form of invariance under a group of transformations, dating back at least to the work of Hoeffding (1952), which generalized older ideas based on testing permutation-invariance (Fisher, 1935; Pitman, 1937). More recent developments in the area pertain to randomized versions of tests for invariance (Romano, 1988, 1989; Hemerik and Goeman, 2017; Hemerik and Goeman, 2018; Dobriban, 2022; Ramdas et al., 2023; Koning and Hemerik, 2024) or approximate invariance (Canay, Romano and Shaikh, 2017); see Ritzwoller, Romano and Shaikh (2024) for a recent review. One takeaway from that line of work is that group-based randomization tests have finite-sample Type I error control (Hoeffding, 1952), power against alternatives (Romano, 1989; Dobriban, 2022), and achieve various notions of optimality and robustness (Romano, 1990; Kashlak, 2022; Koning and Hemerik, 2024).
The literature on group-based randomization tests deals with the invariance of an unknown probability distribution $P_{X}$ from which samples $X_{1},\dotsc,X_{n}$ are obtained. In this paper, we initiate the study of hypothesis tests for the symmetry of a conditional distribution, $P_{Y|X}$, to be tested from samples $(X_{1},Y_{1}),\dotsc,$ $(X_{n},Y_{n})$. Under certain conditions, this can be reduced to a problem for which group-invariance tests are applicable, but those conditions are not satisfied in many settings. The primary motivation for conditional tests is the central role played by symmetry in modern science (Rosen, 1995; Gross, 1996) and the increasingly data-driven nature of scientific discovery, particularly in settings in which the relationships between multiple objects are a key feature (Kofler and Schlötterer, 2012; ATLAS Collaboration, 2017; Collins, Howe and Nachman, 2018; Cristadoro, Degli Esposti and Altmann, 2018; Chmiela et al., 2018; Karagiorgi et al., 2022; Birman et al., 2022; Yim et al., 2023; Watson et al., 2023). A secondary motivation is the prominence of symmetry in certain prediction problems from machine learning (Bronstein et al., 2021).
Figure 1: First column: a random sample from a marginal distribution on $\mathbb{R}^{2}$. A marginal distribution is invariant with respect to rotations about the origin if rotating the distribution does not change it. Second and third columns: conditional distributions $P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ for $Y\in\mathbb{R}^{2}$ (triangle) given three different values of $x\in\mathbb{R}^{2}$ (circle), matched by color. A conditional distribution is rotationally-equivariant if rotations of $x$ transform the distribution of $Y|X=x$ with a corresponding rotation, and conditionally invariant if rotations of $x$ do not change the distribution of $Y|X=x$.
Figure 1 illustrates the differences between marginal and conditional symmetries with a simple example. Whereas symmetry of a probability distribution manifests only as invariance under a group of transformations, a conditional distribution may exhibit a form of symmetry that only appears in relationships between variables, known as equivariance: a transformation of $X$ induces a corresponding transformation of $Y$. As a consequence, testing for conditional symmetry presents certain challenges absent in tests for (marginal) distributional symmetry. In the latter, all tests can be understood as variations on averaging over the group with respect to the “uniform distribution”, a property that stems from the fact that every invariant distribution can be factored into the product of the group’s uniform distribution and a distribution corresponding to an independent random variable. Conditional symmetries do not have such a neat and convenient characterization, in part because they may hold with respect to a non-compact group, for which a uniform distribution does not exist. This is in contrast to distributional invariance, which may hold only with respect to a compact group. Beyond the technical challenges of non-compact groups, there are additional statistical and computational challenges to overcome.
One contribution of the present work is a number of characterizations of conditional symmetry that can be used as the basis for a hypothesis test. The first, which generalizes in practically important ways a result of Bloem-Reddy and Teh (2020), characterizes conditional symmetry as a certain conditional independence relationship. However, generic tests for conditional independence that control Type I error over the entire class of conditionally independent distributions are known to have poor statistical properties (Shah and Peters, 2020). To that end, we establish further characterizations of conditional symmetry that can be used as the basis for group-based conditional randomization tests (Candes et al., 2018), which restrict the null hypothesis class by conditioning on part of the data and therefore have much better statistical properties than the aforementioned conditional independence tests.
Related work.
To our knowledge, the existing literature does not contain any general-purpose tests for conditional symmetry. In addition to typical hypothesis testing applications, such a tool would also be of interest as model-checking criteria in prediction applications where, for example, there is reason to believe that transforming the input in a particular way is equivalent to a corresponding transformation in the output. Prediction problems with symmetry have been a subject of intense study in machine learning; Bronstein et al. (2021) provide a recent review. Besides the literature on group-based randomization tests, some recent methods in machine learning aim to directly learn or estimate symmetry groups from data (Krippendorf and Syvaeri, 2020; Zhou, Knowles and Finn, 2021; Desai, Nachman and Thaler, 2022; Dehmamy et al., 2021; Yang et al., 2023). They do not address the corresponding inferential questions, and the present work can be viewed as a complement to those estimation methods.
To the best of our knowledge, the only work that addresses any notion of testing for conditional symmetry is that of Christie and Aston (2023), who proposed tests for the invariance of the conditional expectation $f(x)=\mathbb{E}[Y|X=x]$, $f\colon\mathbf{X}\to\mathbb{R}$, under the action of some compact group. [^1] We note that the primary aim of Christie and Aston (2023) is to estimate the largest group under which $f$ is invariant, which amounts to performing a collection of tests over a subgroup lattice of some candidate largest group. In principle, our tests could be substituted into their procedure, though we do not address that problem here. Those authors’ focus on testing for the invariance of a function leads them in a different direction, and their tests are limited in their applicability compared to the tests proposed here.
Contributions.
The present work formulates a general framework for testing the invariance or equivariance of a conditional distribution under the action of a specified locally compact group $\mathbf{G}$, based on a single independent and identically distributed (i.i.d.) sample. These tests can be used to detect the presence or absence of a particular conditional symmetry in data, with that symmetry specified by hypothesis or by a data-driven method that has learned the symmetry. In situations with known or assumed symmetry, these tests may also be used as model-checking criteria for models meant to exhibit that symmetry.
In addition to a general formulation of symmetry-based conditional randomization tests for which we prove Type I error control properties, we provide specific kernel-based instantiations of our tests, which allow these tests to be used on any data type for which a kernel can be defined. Using a test statistic based on kernel maximum mean discrepancy, we derive a finite-sample lower bound on the test’s power against non-symmetric alternatives, which also implies that the test is consistent as the number of observations $n\to\infty$.
We demonstrate the application of our tests on synthetic data, and on two problems from high-energy particle physics. We also demonstrate the tests’ use for measuring the symmetry properties of a MNIST classifier that is trained to be invariant by using data augmentation. Computer code required to run the experiments can be found on an associated GitHub repository. [^2] https://github.com/chiukenny/Tests-for-Conditional-Symmetry
Outline.
The paper is structured as follows. In Section 2, we review the necessary basics of groups (additional technical aspects are contained in Section A.1), and give an overview of randomization tests for group invariance. We also fill a gap in the literature on such tests, showing that they are conditional level-$\alpha$ tests for any compact group that acts continuously (Theorem 2.2). In Section 3, we introduce the framework for conditional symmetry and describe some examples. Section 4 is the main technical section, in which conditional symmetry is characterized via conditional independence and certain distributional identities. We propose a general framework for conditional randomization tests of symmetry based on those results in Section 5. In Section 6, we describe specific nonparametric tests based on kernel maximum mean discrepancy, along with finite-sample lower bounds on power against non-symmetric alternatives. Section 7 contains experiments on synthetic data and on particle physics data, as well as a demonstration of how the tests can be used to assess goodness-of-fit for flexible prediction models meant to exhibit symmetry.
2 Background
Before reviewing the basics of groups and randomization tests for group invariance, we briefly introduce some technical necessities and notation. Throughout, we use $\mathbf{X}$ to denote a topological space and $\mathbf{S}_{\mathbf{X}}$ its Borel $\sigma$-algebra, so that $(\mathbf{X},\mathbf{S}_{\mathbf{X}})$ is a standard Borel measurable space. Unless stated otherwise, we refer to $\mathbf{X}$ as a measurable space, and likewise for $\mathbf{Y}$ and $\mathbf{M}$. Let $\mathcal{P}(\mathbf{X})$ denote the set of all probability measures on $\mathbf{X}$. For $P\in\mathcal{P}(\mathbf{X})$, we write $X\sim P$ for a random variable $X$ sampled from $P$. A vector or tuple of random variables is denoted by $X_{1:n}\coloneq(X_{1},\dotsc,X_{n})$. For any measurable function $f$ on $\mathbf{X}$, let $f_{*}P$ denote the pushforward measure with $f_{*}P(A)=P(f^{-1}(A))$ for all measurable sets $A\in\mathbf{S}_{\mathbf{X}}$. We use $\delta_{x}$ to denote the Dirac measure at a point $x$, and =$\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}$ to denote equality in distribution. Let $[n]$ denote the set $\{1,2,\dotsc,n\}$.
2.1 Groups and group actions
This work focuses on equivariant symmetries that can be encoded in the language of group theory, which comprise many of the symmetries of interest in modern statistical, computational, and scientific problems. We assume a basic familiarity with groups and group actions; see Appendix A for background on groups. A thorough treatment of group theory and its applications in statistical contexts can be found in (Eaton, 1989; Wijsman, 1990).
We denote a group by $\mathbf{G}$ and its elements by $g$. For convenience, group composition is written multiplicatively, i.e, $g_{1}g_{2}=g_{1}\cdot g_{2}$. We assume that $\mathbf{G}$ has a topology that is locally compact, second countable, and Hausdorff (lcscH), and which makes the group operations continuous. We may then take $\mathbf{S}_{\mathbf{G}}$ as the Borel $\sigma$-algebra, making $\mathbf{G}$ a standard Borel space. For $A\subseteq\mathbf{G}$ and $g\in\mathbf{G}$, we write $gA=\{gh:h\in A\}$ and $Ag=\{hg:h\in A\}$. A measure $\nu$ on $\mathbf{G}$ is said to be left-invariant if $\nu(gA)=\nu(A)$ for all $A\in\mathbf{S}_{\mathbf{G}}$, and right-invariant if $\nu(Ag)=\nu(A)$. When $\mathbf{G}$ is lcscH, there exist left- and right-invariant $\sigma$-finite measures $\lambda_{\mathbf{G}}$ and $\tilde{\lambda}_{\mathbf{G}}$, known as left- and right-Haar measures, respectively, that are unique up to scaling (Folland, 2016, Ch. 2.2). We use $\lambda$ to denote left-Haar measure. If $\mathbf{G}$ is compact, then $\lambda=\tilde{\lambda}$, and the unique normalized Haar measure acts as the uniform probability measure over the group.
A group $\mathbf{G}$ acts measurably on a set $\mathbf{X}$ if the group action $(g,x)\mapsto gx$ is a measurable function. Throughout, all actions are assumed to be continuous (and therefore measurable), and for convenience we simply say that $\mathbf{G}$ acts on $\mathbf{X}$. For a set $A\subseteq\mathbf{X}$, $gA=\{gx:x\in A\}$. The stabilizer subgroup of $x\in\mathbf{X}$ is the subset of group actions that leave $x$ invariant, denoted $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$. The action is free (or exact) if $gx=x$ implies that $g=\textrm{id}$, in which case $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$. The $\mathbf{G}$-orbit of $x\in\mathbf{X}$ is the subset of all points in $\mathbf{X}$ reachable through a group action in $\mathbf{G}$, denoted $O(x)=\{gx:g\in\mathbf{G}\}$. Each orbit is an equivalence class of $\mathbf{X}$, where two points are equivalent if and only if they belong to the same orbit. If $\mathbf{X}$ has only one orbit, then the action is transitive.
In the analysis of probabilistic aspects of group actions, measurability issues can arise without regularity conditions. The key regularity condition that we assume in this work is that the group action is proper. Properness is a common assumption in statistical applications of group theory (e.g., Eaton, 1989; Wijsman, 1990; McCormack and Hoff, 2023) and is satisfied in many settings of interest. We discuss the implications of this assumption in Section A.1.
2.2 Testing for distributional invariance
We briefly review the basic framework and properties of tests for group invariance of a probability measure. A function $f$ with domain $\mathbf{X}$ is $\mathbf{G}$-invariant if $f(gx)=f(x)$, all $x\in\mathbf{X},g\in\mathbf{G}$. This specializes to probability measures.
Definition 2.1.
A probability measure $P$ on $\mathbf{X}$ is $\mathbf{G}$-invariant if $g_{*}P(A)\coloneq P(g^{-1}A)=P(A)$ for all $g\in\mathbf{G}$, $A\in\mathbf{S}_{\mathbf{X}}$.
For a specified finite or compact group $\mathbf{G}$, consider the hypotheses
[…]
Following Koning and Hemerik (2024), a level-$\alpha$ test based on statistic $T$ is
$\phi_{\alpha}(X)=\mathbbm{1}\{T(X)>q_{T(GX)}(1-\alpha)\}\;,$ (2)
where $q_{T(GX)}(1-\alpha)$ is the quantile function of $T(GX)$ at $1-\alpha$, with $X\sim P$ and $G\sim\lambda$, a random element of $\mathbf{G}$ sampled from Haar measure on $\mathbf{G}$. Showing that this test has level $\alpha$ is straightforward, using the fact that $P$ is $\mathbf{G}$-invariant if and only if $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GX$. Hence $\mathbb{E}[\phi_{\alpha}(X)]=\mathbb{E}[\phi_{\alpha}(GX)]\leq\alpha$.
Algorithm 1 Randomization test for $\mathbf{G}$-invariance
1:inputs: $X_{1:n}$, $B$, $T$, $\alpha$
2:compute $T(X_{1:n})$
3:for $b$ in $1,\dotsc,B$ do
4:     sample $G_{i}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}\lambda$
5:     set $X_{1:n}^{(b)}\coloneqq(G_{1}^{(b)}X_{1},\dotsc,G_{n}^{(b)}X_{n})$
6:     compute $T(X_{1:n}^{(b)})$
7:end for
8:compute $p$-value $p_{B}$ as
$p_{B}\coloneqq\frac{1+\sum_{b=1}^{B}\mathbbm{1}\{T(X_{1:n}^{(b)})\geq T(X_{1:n})\}}{1+B}$ (3)
9:reject $H_{0}$ if $p_{B}\leq\alpha$
For small and finite $\mathbf{G}$, the quantile $q_{T(GX)}(1-\alpha)$ can be computed exactly by applying each element of $\mathbf{G}$ to $X$. However, in the case that $\mathbf{G}$ is large or uncountable, randomization tests may be used. Algorithm 1 shows the basic recipe. We note that Algorithm 1 differs slightly from the conventional procedure described in the literature (e.g., Hemerik and Goeman, 2018), in which a randomization is performed by sampling a single transformation $G^{(b)}\sim\lambda$ and applying that to the entire sample $X_{1:n}$ (as a single multivariate observation). Our procedure is closer to that described by Romano (1989) and can be seen as a generalization of the conventional procedure to (possibly multivariate) i.i.d. observations. We recover the conventional procedure by treating $X_{1:n}$ as a single observation in the space $\mathbf{X}^{n}$ and taking the group action $G^{(b)}=G^{(b)}_{1}\times\dotsc\times G^{(b)}_{1}$ as an action on $\mathbf{X}^{n}$.
For the conventional procedure, an “unconditional” proof that the resulting test is level-$\alpha$ was given for discrete $\mathbf{G}$ by Hemerik and Goeman (2017). Their proof method can also be used without issue for (uncountable) compact $\mathbf{G}$. A second proof, in which the level of the test holds conditionally on the orbit of $X$, was given by Hemerik and Goeman (2018). That conditional proof, however, relies on the countability of the orbit, which fails when $\mathbf{G}$ is uncountable (e.g., the group of continuous rotations). Theorem 2.2 below fills this gap in the literature, holding for randomization tests of invariance for compact groups, conditioned on the orbit. Our proof relies on a sufficiency argument, thereby exposing randomization tests for invariance to be co-sufficient sampling methods (Barber and Janson, 2022). The high-level idea is as follows. A maximal invariant is a $\mathbf{G}$-invariant function, $M\colon\mathbf{X}\to\mathbf{M}$, that takes a different value on each orbit, so that $M(x)=M(x^{\prime})$ if and only if $gx=x^{\prime}$ for some $g\in\mathbf{G}$. It is known that any maximal invariant is a sufficient statistic for $\mathcal{P}^{\circ}(\mathbf{X})$, the set of $\mathbf{G}$-invariant probability distributions on $\mathbf{X}$ (Farrell, 1962; Dawid, 1985; Bloem-Reddy and Teh, 2020). Because a maximal invariant $M(X)$ uniquely identifies an orbit $O(X)$, the conditional distribution of $X$ given $M(X)$ is the distribution over $O(X)$ induced by pushing $\lambda$ (the uniform measure on $\mathbf{G}$) through the group action on that orbit. This holds for every invariant probability measure $P\in\mathcal{P}^{\circ}(\mathbf{X})$, even when the action of $\mathbf{G}$ on $\mathbf{X}$ has non-trivial stabilizers. Hence, conditioning on the observed maximal invariants and randomizing with $G\sim\lambda$ generates a set of conditionally i.i.d. samples, and the usual analysis of (conditional) Monte Carlo tests can be applied. Moreover, because every distribution in the null hypothesis set shares the same conditional distribution, $p_{B}$ is approximately uniformly distributed, with $p_{B}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}U\sim\textrm{Unif}[0,1]$ as $B\to\infty$.
Maximal invariants are not unique; we use $M$ to denote a generic maximal invariant, and $M(X)_{1:n}$ to denote the element-wise maximal invariants of $X_{1:n}$.
Theorem 2.2.
Let $\mathbf{G}$ be a compact group that acts on $\mathbf{X}$. Let $M$ be any maximal invariant. Assume that $X_{1:n}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P$. For test statistic $T$ and randomization test in Algorithm 1, asume that $\mathbb{P}\left(T(X^{(b)}_{1:n})=T(X^{(b^{\prime})}_{1:n})\right)=0$ for $b\neq b^{\prime}$. For any fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$, if $P$ is $\mathbf{G}$-invariant then
[…]
The proof of Theorem 2.2 is in Section B.1. As noted by Dufour and Neves (2019), if $\alpha(B+1)$ is an integer, then the inequality in (4) becomes equality. If the probability of ties is non-zero (i.e., if $T$ is supported on a discrete set), then a modified version with randomized tie-breaking yields a level-$\alpha$ test (Dufour, 2006; Hemerik and Goeman, 2018).
An important practical observation is that all of the above holds even for (possibly randomized) test statistics that depend on the observed data through $M(X)_{1:n}$. That means that functional parameters for $T$ can be selected in a data-driven way, as long as the procedure only uses $M(X)_{1:n}$. For example, the power of nonparametric tests based on kernel maximum mean discrepancy are known to depend greatly on the kernel function parameters (Ramdas et al., 2015), so our implementation relies on selecting parameters using the observed $M(X)_{1:n}$. This extends and further justifies an idea introduced by Biggs, Schrab and Gretton (2024) in the context of nonparametric permutation tests, which they allowed to depend on permutation-invariant functions of the data.
Example 1 (Test of 2D rotational invariance).
Consider a setting similar to the one visualized in the first column of Figure 1, where $X$ is a random vector in $\mathbb{R}^{2}$. Suppose that $X=ZR_{\theta}$, where $Z$ is a 2D Gaussian random variable centered on the point $(1,0)$ with identity covariance and $R_{\theta}$ is a random 2D rotation matrix of angle $\theta$,
$R_{\theta}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\;.$ (5)
If $\theta$ is sampled over some range, it is straightforward to see that the marginal distribution of $X$ is invariant with respect to $\mathrm{SO}(2)$—the group of 2D rotations about the origin $(0,0)$—if and only if $\theta$ is uniform on $[0,2\pi)$.
Figure 2: Test rejection rates for the distribution described in Example 1 at sample size $n=100$ when $\theta$ is sampled uniformly from $[0,p)$. The distribution is $\mathrm{SO}(2)$-invariant only when $p=2\pi$. Test statistics (FS and SK) are described in Section 6; RN indicates randomization using Algorithm 1 and BS indicates a two-sample permutation test between $X_{1:n}$ and a single randomization generated as in Algorithm 1.
Using the kernel-based test statistics described in Section 6 within the randomization procedure in Algorithm 1, Figure 2 (RN-FS and RN-SK) shows the rejection rates, estimated over $1000$ simulations, of several tests for rotational invariance on samples $X_{1:n}$, where $\theta_{1},\dotsc,\theta_{n}$ are sampled i.i.d. over $[0,p)$ with $p\in\{\nicefrac{{\pi}}{{2}},\pi,\nicefrac{{3\pi}}{{2}},2\pi\}$. As $p\to 2\pi$, the distribution of $X$ becomes increasingly more invariant, which is reflected in the power of the tests decreasing with $p$. When $p=2\pi$, the tests reject at a rate of approximately $\alpha=0.05$. As a baseline, Figure 2 (BS-FS and BS-SK) also shows the results of a baseline test: comparing $X_{1:n}$ with a single randomization set $X_{1:n}^{(0)}$ using a two-sample permutation test. Histograms showing the $p$-values for these tests can be found in Section D.1.
3 Symmetry of conditional distributions
The primary focus of this paper is on a different notion of probabilistic symmetry, pertaining to conditional distributions. Here, we review the basic concepts needed to describe the problem. Suppose $\mathbf{G}$ acts on $\mathbf{X}$ and also on another set $\mathbf{Y}$; the group action may be different on each. A function $f\colon\mathbf{X}\to\mathbf{Y}$ is $\mathbf{G}$-equivariant if $f(gx)=gf(x)$, for all $x\in\mathbf{X},\ g\in\mathbf{G}$. In the case that $\mathbf{G}$ acts trivially on $\mathbf{Y}$, equivariance specializes to invariance.
We say that $P_{X,Y}$ is jointly $\mathbf{G}$-invariant if it is invariant in the sense of Definition 2.1 extended to $\mathbf{G}$ acting on $\mathbf{X}\times\mathbf{Y}$. In addition to joint invariance, we may define symmetry for a conditional distribution. We write the disintegration of $P_{X,Y}$ as $P_{X,Y}=P_{X}\otimes P_{Y|X}$, where $P_{Y|X}$ denotes a regular conditional probability of $Y$ given $X$ (i.e., a Markov probability kernel from $\mathbf{X}$ to $\mathbf{Y}$).
Definition 3.1.
The conditional distribution $P_{Y|X}$ is $\mathbf{G}$-equivariant if
$P_{Y|X}(B\mid x)=P_{Y|X}(gB\mid gx)\;,\quad x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G}\;.$ (6)
$P_{Y|X}$ is $\mathbf{G}$-invariant if
$P_{Y|X}(B\mid gx)=P_{Y|X}(B\mid x)\;,\quad x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G}\;.$ (7)
As a simple example of equivariance, consider the case in which $Y\in\mathbb{R}^{d}$ is conditionally Gaussian-distributed and is acted on by $\text{SO}(d)$, the group of $d$-dimensional rotations.
Example 2.
For random vectors $X$, $Y$ in $\mathbb{R}^{d}$, suppose that $Y|X\sim\mathsf{N}(X,\mathbf{I}_{d})$. It is straightforward to check that $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant. If $\varepsilon\sim\mathsf{N}(0,\mathbf{I}_{d})$ with $\varepsilon{\;\perp\!\!\!\perp\;}X$ so that $Y=X+\varepsilon$, then for any $g\in\mathrm{SO}(d)$, $A\subseteq\mathbb{R}^{d}$,
$P_{Y|X}(gY\in A\mid X)$ $=P_{Y|X}(gX+g\varepsilon\in A\mid X)$
$=P_{Y|X}(gX+\varepsilon\in A\mid X)=P_{Y|X}(Y\in A\mid gX)\;.$
If, instead, $Y|X\sim\mathsf{N}(\|X\|,\Sigma)$, then $P_{Y|X}$ is $\text{SO}(d)$-invariant.
We use the term “$\mathbf{G}$-symmetric” as a catch-all for either of the conditional symmetries in Definition 3.1. Similar to the hypotheses (1) in testing distributional invariance, the primary aim of this paper is to test the hypotheses
$\begin{split}H_{0}\colon P_{Y|X}\text{ is $\mathbf{G}$-symmetric}\qquad\text{versus}\qquad H_{1}\colon P_{Y|X}\text{ is not $\mathbf{G}$-symmetric.}\end{split}$ (8)
3.1 Representatives and inversions
$[\mathbf{X}]$$(0,0)$$O(x)$$x=\tau(x)[x]$$[x]=\rho(x)$$\tau(x)$$M(x)$
Figure 3: A graphical depiction of a point $x\in\mathbb{R}^{2}$, its orbit $O(x)=\{x^{\prime}\in\mathbb{R}^{2}:\|x^{\prime}\|_{2}=M(x)\}$, and its orbit representative $[x]\in[\mathbf{X}]=\{(z,0):z\geq 0\}$, and the role of the orbit selector $\rho$, maximal invariant $M$, and representative inversion $\tau$ in the context of two-dimensional rotations about the origin $(0,0)$. Note that $[x]$ itself is also a maximal invariant.
One property distinguishing marginal invariance from conditional symmetry is that marginal invariance can only hold with respect to compact groups, whereas the group can be non-compact in the case of conditional symmetries. In principle, that makes conditional symmetries applicable to a wider range of problems, though it also makes them more challenging to formulate and analyze. In order to address those challenges, we rely on special structure arising from group actions, beyond what is typically needed for tests of distributional invariance. For technical details on measurability issues that can arise, the interested reader may refer to Section A.1. Throughout, we assume that the group action is proper so that conditions needed for measurability are satisfied.
On each orbit $O(x)$, we can choose an element $[x]\in O(x)$ as the orbit representative, so that $[x]=gx$ for some $g\in\mathbf{G}$. $[x]$ can be thought of as an “origin” on the orbit. The properties that we describe below do not depend on which element is chosen as the representative, but the underlying structure of the properties will be relative to that choice. For a particular choice of representatives, the subset of $\mathbf{X}$ consisting of each orbit’s representative is denoted by $[\mathbf{X}]$. Note that $[\mathbf{X}]\cap O(x)=\{[x]\}$. A function $\rho\colon\mathbf{X}\to[\mathbf{X}]$ that maps elements of $\mathbf{X}$ onto their corresponding orbit representatives is called an orbit selector. Any orbit selector is a maximal invariant by definition. Figures 3 and 4 depict these quantities for two- and three-dimensional rotations, respectively.
$O(x)$$[\mathbf{X}]$$(0,0,0)$$[x]=\rho(x)$$x$$g_{1}$$g_{2}$
Figure 4: A graphical depiction of a point $x\in\mathbb{R}^{3}$, its orbit $O(x)$ and orbit representative $[x]\in[\mathbf{X}]$ in the context of three-dimensional rotations about the origin $(0,0,0)$. Depicted are two rotations $g_{1}$ and $g_{2}$ that take $[x]$ to $x$. The inversion kernel $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ is the uniform distribution over $\{g\in\mathrm{SO}(3):x=g[x]\}=g\mathbf{G}_{[x]}$.
Recall that the stabilizer subgroup of $x$ is $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$, and that the action is free if $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$. A $\mathbf{G}$-equivariant function $\tau\colon\mathbf{X}\to\mathbf{G}$ is called a representative inversion if $\tau(x)\rho(x)=x$ for all $x\in\mathbf{X}$. As the name suggests, $\tau$ returns the element of $\mathbf{G}$ that moves, or “inverts”, the representative $\rho(x)$ to $x$; the inverse group element $\tau(x)^{-1}$ moves $x$ to $\rho(x)$. In order for $\tau$ to be uniquely defined, the group action must be free. If it is not, an equivariant inversion probability kernel (inversion kernel for short) $P_{\tilde{\tau}|X}\colon\mathbf{X}\times\mathbf{S}_{\mathbf{G}}\to[0,1]$ can be used in place of $\tau$. If $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$, then $x\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{=}}\tilde{\tau}\rho(x)$. At a high level, one may think of the inversion kernel $P_{\tilde{\tau}|X}$ as the uniform distribution on the left coset $g\mathbf{G}_{\rho(x)}$, where $g\rho(x)=x$. Assuming that the group action is proper (see Section A.1) ensures that the left coset is compact and a uniform distribution is well-defined. When the action is free, the inversion kernel simplifies to $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)=\delta_{\tau(x)}({\,\vbox{\hbox{\tiny$\bullet$}}\,})$. Figures 3 and 4 depict these structures for $\text{SO}(2)$ and $\text{SO}(3)$ acting on $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$, respectively. The main difference is that the stabilizer subgroup of $x\in\mathbb{R}^{2}$ is trivial while that of $x\in\mathbb{R}^{3}$ is not.
In the setting of distributional invariance in Section 2.2, the conditional level stems from the decomposition $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)$ with $G{\;\perp\!\!\!\perp\;}X$ and $G\sim\lambda$ (see Section B.1). To test for conditional symmetry, we will need a similar decomposition with a random element $\Gamma\in\mathbf{G}$, so that $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}\Gamma\rho(X)$. However, here the conditional distribution over $\mathbf{G}$ given an orbit of $X$ will in general depend on the orbit and not be invariant. We use $\Gamma$ to denote the random variable with this conditional distribution, which can be written as
$P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid[x])=\int_{\mathbf{X}}P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid x^{\prime})P_{X|[X]}(dx^{\prime}\mid[x])\;,$ (9)
where $P_{X|[X]}$ is the conditional distribution of $X$ given its orbit. We refer to $P_{\Gamma|[X]}$ as the group-orbit kernel as it is the distribution on $\mathbf{G}$ induced by restricting $P_{X}$ to the orbit $O([x])$. The notation indicates that it is conditioned on a particular orbit represented by $[x]$, but it can be induced by conditioning on any maximal invariant. The group-orbit kernel has the property that if $\Gamma\sim P_{\Gamma|[X]}$, then $(\rho(X),X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X))$. (See Lemma A.2 for a proof.)
In some cases, a representative inversion can still be (non-uniquely) defined when the action is not free (see Example 3), in which case an equivalent inversion kernel can be defined as $P_{\tilde{\tau}|X}^{\prime}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)\coloneqq P_{\tilde{\tau}|X}(\tau(x)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid[x])$. Inversion kernels do not seem to have been used previously in statistics or machine learning, likely owing to their relatively recent appearance in probability (Kallenberg, 2011). However, the deterministic representative inversions have been seen in the recent machine learning literature (Bloem-Reddy and Teh, 2020; Winter et al., 2022; Kaba et al., 2023).
As previously mentioned, when $\mathbf{G}$ is non-compact, there may not exist measurable maximal invariants, orbit selectors, and inversion kernels. We assume that the action of $\mathbf{G}$ on any set is continuous and proper; these conditions are implicit in statements such as “let $\mathbf{G}$ be a group that acts on $\mathbf{X}$”. For our purposes, properness is sufficient for the existence of measurable versions of these objects. See Section A.1 for more details.
The following examples illustrate the main ideas, and provide background for our experiments in Section 7.
Example 3.
Let $X$ be a random vector in $\mathbb{R}^{d}$, acted on by $\text{SO}(d)$. A convenient maximal invariant is $M(X)=\|X\|$. The set of orbit representatives can be chosen to be the points on the axis with unit basis vector $\mathbf{e}_{1}=[1,0,\dotsc,0]^{\top}$. Then for each $x\in\mathbb{R}^{d}$, $\rho(x)=\|x\|\mathbf{e}_{1}$. For $d=2$, the action is free; for $d>2$, the stabilizer subgroup $\mathbf{G}_{\rho(x)}$ is the set of $d$-dimensional rotations around the axis corresponding to $\mathbf{e}_{1}$. One may construct a representative inversion function corresponding to $\rho(x)=\|x\|\mathbf{e}_{1}$ by, for example, rotating $\|x\|\mathbf{e}_{1}$ to $x$ in the 2D subspace spanned by $\nicefrac{{x}}{{\|x\|}}$ and $\mathbf{e}_{1}$. That is, let $\tilde{x}\coloneqq\nicefrac{{(x-\langle\mathbf{e}_{1},x\rangle\mathbf{e}_{1})}}{{\|x-\langle\mathbf{e}_{1},x\rangle\mathbf{e}_{1}\|}}$, so that $[\mathbf{e}_{1},\ \tilde{x}]$ is a matrix in $\mathbb{R}^{d\times 2}$ whose columns form an orthonormal basis for the 2D subspace spanned by $\nicefrac{{x}}{{\|x\|}}$ and $\mathbf{e}_{1}$. Now let $\theta_{x}$ be such that $\cos(\theta_{x})=\langle\mathbf{e}_{1},\nicefrac{{x}}{{\|x\|}}\rangle$, and $R_{\theta}$ the standard 2D rotation matrix of angle $\theta$ as defined in (5). Then the $d$-dimensional rotation defined by
$\tau(x)=\mathbf{I}_{d}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top}-\tilde{x}\tilde{x}^{\top}+[\mathbf{e}_{1},\ \tilde{x}]R_{\theta_{x}}[\mathbf{e}_{1},\ \tilde{x}]^{\top}$ (10)
satisfies $\tau(x)\cdot\|x\|\mathbf{e}_{1}=x$. A sample from the corresponding inversion kernel is then generated by taking a uniform random $(d-1)$-dimensional rotation $H$ and extending it to a $d$-dimensional rotation $H^{\prime}$ that fixes $\mathbf{e}_{1}$, so that $\tau(x)H^{\prime}$ has distribution $P_{\tilde{\tau}|X}$.
Example 4.
The Lorentz group, or indefinite orthogonal group $\text{O}(1,3)$, is a fundamental symmetry group in high-energy physics, and is central to the physical description of particles moving in laboratories at velocities approaching the speed of light (Bogatskiy et al., 2020). We give only the relevant details here, largely following (Bogatskiy et al., 2020); more details can be found in standard textbook references (e.g., Sexl and Urbantke, 2001; Cottingham and Greenwood, 2007). $\text{O}(1,3)$ is the group of linear isometries of Minkowski spacetime, on which the theory of special relativity is constructed. For a particle’s four-momentum, $x=(E,p_{1},p_{2},p_{3})\in\mathbb{R}^{4}$, the group preserves the quadratic form
$Q(x)=E^{2}-p_{1}^{2}-p_{2}^{2}-p_{3}^{2}\;,$ (11)
and so by definition, $Q$ is a maximal invariant. The subgroup that preserves the orientation of time and space (hence viewed as corresponding to physically realistic transformations) is known as the restricted Lorentz group, denoted $\mathrm{SO}^{+}(1,3)$. Mathematically, the restricted Lorentz group is a non-compact, non-commutative, connected real Lie group (a differentiable manifold) that consists of transformations $x\mapsto g_{R}g_{\beta}x$ composed of boosts $g_{\beta}$ and three-dimensional rotations $g_{R}$ that act on the directional entries $(p_{1},p_{2},p_{3})$. For $\beta\in(-1,1)$ and $R\in\textrm{SO}(3)$, transformations in $\mathrm{SO}^{+}(1,3)$ of a particle’s four-momentum $x$ are encoded by
$g_{\beta}\begin{bmatrix}E\\ p_{1}\\ p_{2}\\ p_{3}\end{bmatrix}$ $=\begin{bmatrix}(1-\beta^{2})^{-\frac{1}{2}}(E-\beta p_{1})\\ (1-\beta^{2})^{-\frac{1}{2}}(-\beta E+p_{1})\\ p_{2}\\ p_{3}\end{bmatrix}\;,$ $g_{R}x$ $=\begin{bmatrix}1&0&0&0\\ 0&R_{11}&R_{12}&R_{13}\\ 0&R_{21}&R_{22}&R_{23}\\ 0&R_{31}&R_{32}&R_{33}\end{bmatrix}x\;.$
These transformations are sometimes referred to as “hyperbolic rotations” that give rise to orbits whose one-dimensional projections are hyperbolas. The set of orbit representatives can be chosen as $\rho(x)=(E_{\rho(x)},1,0,0)$, so that $[\mathbf{X}]=\{(E_{\rho},1,0,0):E_{\rho}\in\mathbb{R}_{+}\}$, with $E_{\rho(x)}^{2}-1=Q(x)$. For any $x=(E,p_{1},p_{2},p_{3})$, a sample from the inversion kernel $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ can be obtained in two steps: first, calculate $\beta=\frac{E_{\rho(x)}-E\sqrt{p_{1}^{2}+p_{2}^{2}+p_{3}^{2}}}{1+E^{2}}$; second, sample a 3D rotation $R$ uniformly at random from $\mathbf{G}_{(1,0,0)}$ and compute $\tau((p_{1},p_{2},p_{3}))$ as in (10). Then $x=g_{\tau((p_{1},p_{2},p_{3}))}G_{R}g_{\beta}\rho(x)$.
4 Characterizations of conditional symmetry
Definition 3.1 does not provide an actionable characterization of conditional symmetry with which to build hypothesis tests. In this section, we establish a number of characterizations of conditional symmetry. While they have immediate implications for testing, a detailed discussion on the construction of these tests will be the main focus of Section 5. The first characterization is in terms of a particular form of conditional independence that elucidates the structure of equivariant conditional distributions.
Equivariant Markov kernels arise naturally in the disintegration of jointly invariant measures. Under general conditions, a measure is jointly invariant if and only if it disintegrates into an invariant marginal measure and an equivariant Markov kernel. A measure-theoretic proof is provided by Kallenberg (2017, Theorem 7.6). Bloem-Reddy and Teh (2020) connected this to conditional independence, showing that for compact $\mathbf{G}$ and $\mathbf{G}$-invariant $P_{X}$ such that a measurable representative inversion $\tau$ exists, $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if
$X{\;\perp\!\!\!\perp\;}\tau(X)^{-1}Y\mid M(X)\;.$ (12)
Informally, the relation in (12) says that given knowledge of the orbit through a maximal invariant $M(X)$, the distribution of the transformed variable $\tau(X)^{-1}Y$ does not depend on the particular value $X$ takes on its orbit. Figure 5 illustrates this intuition for a simple rotation example: the aligned conditional distributions of $Y|X=x$ for different values of $x$ on the same orbit should be the same distribution; if not, additional dependency between $Y$ and $X$ is indicated.
Figure 5: Three random samples of $Y\in\mathbb{R}^{2}$ (triangle) drawn conditionally on $X\in\mathbb{R}^{2}$ (circle). If the conditional distribution $P_{Y|X}$ is $\mathrm{SO}(2)$-equivariant, then $\tau(x)_{*}P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ is the same for all samples.
The assumptions for (12) in (Bloem-Reddy and Teh, 2020) are somewhat restrictive, as they require $\mathbf{G}$ to be compact and $P_{X}$ to be invariant. In many situations, one or both of those assumptions will not hold. The following theorem substantially generalizes (12), making it applicable to non-compact $\mathbf{G}$ and non-invariant $P_{X}$. Moreover, it does not require the existence of a representative inversion $\tau(x)$.
Theorem 4.1.
Let $\mathbf{G}$ be a lcscH group acting on each of $\mathbf{X}$ and $\mathbf{Y}$, the action on $\mathbf{X}$ proper, with measurable inversion kernel $P_{\tilde{\tau}|X}$. The conditional distribution $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if,
$(\tilde{\tau},X){\;\perp\!\!\!\perp\;}\tilde{\tau}^{-1}Y\mid M(X)\;,\quad\text{with}\quad\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X)\;.$ (13)
The proof can be found in Section B.2. We note the following special cases. If there exists a representative inversion function $\tau\colon\mathbf{X}\to\mathbf{G}$, then condition (13) reduces to (12). If the action of $\mathbf{G}$ on $\mathbf{Y}$ is trivial, then (13) reduces to $X{\;\perp\!\!\!\perp\;}Y\mid M(X)$. Finally, if the action of $\mathbf{G}$ on $\mathbf{X}$ is transitive, then (13) specializes to the (unconditional) independence.
Theorem 4.1 implies that a test for the hypotheses (8) can be performed using a test for the conditional independence in (13). If the action of $\mathbf{G}$ on $\mathbf{X}$ is transitive, then $\mathbf{X}$ has a single orbit and the test simplifies to a test for (unconditional) independence. In principle, any general-purpose test for conditional independence may be used; see Li and Fan (2020) for a recent review of the literature. However, it is known that any test of size $\alpha$ over the entire null hypothesis class has no power against any alternative (Shah and Peters, 2020). The high-level solution is to restrict the null hypothesis class, which is commonly accomplished through problem-specific assumptions (Shah and Peters, 2020; Kim et al., 2022) or modeling some parts of the relevant conditional distributions (Zhang et al., 2011; Pogodin et al., 2024).
One broadly useful set of tools for testing conditional independence are conditional randomization tests (Candes et al., 2018; Berrett et al., 2019; Liu et al., 2022), in which a suitable randomization of the data is used to generate a test that is similar in spirit to Algorithm 1. This motivates a refinement of the characterization of conditional symmetry in Theorem 4.1, on which we will build a conditional randomization test in the next section. The basic idea is that for a pair $(X,Y)$, the “centered” variables $\rho(X)$ and $\tilde{Y}\coloneqq\tilde{\tau}^{-1}Y$ can be used to obtain a randomized pair $(X^{\prime},Y^{\prime})\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(X,Y)$. The randomization involves sampling one or both of
$\tilde{Y}\sim P_{\tilde{Y}|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\rho(X))\;,\quad\Gamma\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\rho(X))\;,$ (14)
with $\tilde{Y}{\;\perp\!\!\!\perp\;}\Gamma\mid\rho(X)$.
Proposition 4.2.
Let $P_{X,Y}=P_{X}\otimes P_{Y|X}$ be a probability measure on $\mathbf{X}\times\mathbf{Y}$. For a lcscH group $\mathbf{G}$ acting on $\mathbf{X}$ and $\mathbf{Y}$, with the action on $\mathbf{X}$ proper, let $P_{\Gamma|[X]}$ be the group-orbit kernel (9). Then the following statements are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E1.
    $(\rho(X),\tilde{Y},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},\Gamma\rho(X),\Gamma\tilde{Y})$, with $\Gamma$ sampled as in (14).
  2. E2.
    $(\rho(X),\tilde{\tau},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{\tau},\tilde{\tau}\rho(X),\tilde{\tau}\tilde{Y})$, with $\tilde{Y}$ sampled as in (14).
  3. E3.
    $(\rho(X),X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X),\Gamma\tilde{Y})$, with $\tilde{Y},\Gamma$ sampled as in (14).
These properties follow as elaborations of Theorem 4.1, and the proof can be found in Section B.2. Taken together, they say that if we are able to sample at least one of $\Gamma$ and $\tilde{Y}$ from the correct distribution conditional on $\rho(X)$, we can test for $\mathbf{G}$-equivariance by ftesting for equality in distribution between the data and the randomization $(X^{\prime},Y^{\prime})\coloneq(\Gamma\rho(X),\Gamma\tilde{Y})$. Crucially, the randomization (and therefore the test) can be conditional on part of the observed data. The part that gets conditioned on are the variable(s) that precede $X,Y$ on the left-hand side of the distributional identity in each property. For example, Property E1 allows us to condition on $(\rho(Y),\tilde{Y})$. In each case, conditioning on part of the data restricts the null hypothesis set to agree with that part of the data, and therefore sidesteps the impossibility result of Shah and Peters (2020).
Proposition 4.2 has two special cases that are important for practical reasons. First, if $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$, the data being used in the test can be reduced to $Y$ only.
Corollary 4.3.
Assume the conditions of Proposition 4.2. If $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$ then the following statements are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E4.
    $(\rho(X),\tilde{Y},Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},\Gamma\tilde{Y})$, with $\Gamma$ sampled as in (14).
  2. E5.
    $(\rho(X),\tilde{\tau},Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{\tau},\tilde{\tau}\tilde{Y})$, with $\tilde{Y}$ sampled as in (14).
  3. E6.
    $(\rho(X),Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\tilde{Y})$, with $\tilde{Y},\Gamma$ sampled as in (14).
The second useful special case occurs when $P_{X}$ is invariant under a compact group, in which case the test for conditional symmetry is equivalent to joint invariance of $P_{X,Y}$. Therefore, in such situations we can test for $\mathbf{G}$-equivariance of $P_{Y|X}$ by testing for joint $\mathbf{G}$-invariance of the distribution $P_{X,Y}$ or by testing for marginal $\mathbf{G}$-invariance of the distribution $P_{Y}$ (when $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$), leveraging the established body of work on group randomization tests for distributional invariance (Section 2.2).
Corollary 4.4.
If $\mathbf{G}$ is compact and $P_{X}$ is $\mathbf{G}$-invariant, the following are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E7.
    $(\rho(X),\tilde{Y},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},G\rho(X),G\tilde{Y})$ with $G\sim\lambda$.
  2. E8.
    $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$, and $Y\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GY\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\tilde{Y}$ for $G\sim\lambda$.
In Section 5, we describe how valid tests for conditional symmetry can be implemented for each of these cases.
Example 5 (Continuation of Example 2).
Consider random vectors $X$ and $Y$ in $\mathbb{R}^{d}$, acted on by $\mathrm{SO}(d)$. As described in Example 3, a convenient maximal invariant is $M(X)=\|X\|$. Then Property E6 says that $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant if and only if $Y\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}\Gamma\tilde{Y}$, where $\tilde{Y}\sim P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\|X\|\mathbf{e}_{1})$, and where $\Gamma\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\|X\|\mathbf{e}_{1})$ is a random $d$-dimensional rotation matrix. Property E5 replaces $\Gamma$ with $\tilde{\tau}$, where $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X)$ (for which the sampling procedure is described in Example 3); alternatively, Property E4 replaces $\tilde{Y}$ with $\tilde{\tau}^{-1}Y$. If $P_{Y|X}=\mathsf{N}(X,\Sigma)$, then $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant if and only if $\Sigma=\mathbf{I}_{d}$.
Example 6 (Continuation of Example 4).
The Standard Model of physics is invariant under the Lorentz group, and violations of $\mathrm{SO}^{+}(1,3)$-invariance could indicate phenomena beyond the Standard Model. (More modestly, violations of $\mathrm{SO}^{+}(1,3)$-invariance could indicate issues with high-energy physics simulators.) As an example, consider the four-momenta of particles produced by high-energy colliders such as the Large Hadron Collider (LHC). High-energy quarks produced in collisions such as those at the LHC quickly decay through a cascading process of gluon emission into collimated sprays of stable hadrons, which are subatomic particles that can be detected and measured (Salam, 2010; Bogatskiy et al., 2020). This spray is known as a jet. According to the Standard Model, the four-momenta of the constituent particles should respect the symmetries of the Lorentz group in the following sense. Let $X$ and $Y$ be the four-momenta relative to the lab frame of any two particles produced by the same collision; hence, they are stochastically dependent. The conditional distribution $P_{Y|X}$ should be $\mathrm{SO}^{+}(1,3)$-equivariant, as a Lorentz transformation of $X$ is equivalent to a change of the lab reference frame, which would then also apply to $Y$. Observe that this is not distributional invariance, because $gX\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\neq}}X$. Example 4 describes how to sample from the inversion kernel, so if samples can be generated from $X^{\prime}\sim P_{X|Q(X)}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|Q(X))$, then $\Gamma=\tilde{\tau}^{\prime}$ with $\tilde{\tau}^{\prime}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X^{\prime})$ and $\tilde{Y}=\tilde{\tau}^{-1}Y$ with $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X)$ satisfy the distributional identities above.
In the experiment in Section 7.4, we test whether this equivariance holds, using the Top Quark Tagging Reference dataset (Kasieczka et al., 2019), which consists of simulated collision events generated by PYTHIA, a Monte Carlo simulator that is widely used in high-energy physics.
5 Randomization tests for conditional group symmetry
The characterizations of conditional symmetry in the previous section suggest a simple way to test for conditional symmetry, in the form of a conditional randomization test (CRT). The basic idea for this setting is as follows. With a set of observations $(X,Y)_{1:n}$, we generate $B$ randomization sets $(X,Y)_{1:n}^{(b)}$ as
$(X,Y)_{i}^{(b)}\coloneq(\mathsf{\Gamma}_{i}^{(b)}\rho(X_{i}),\mathsf{\Gamma}_{i}^{(b)}\mathsf{\tilde{Y}}_{i}^{(b)})\;,\quad b=1,\dotsc,B.$
Proposition 4.2 indicates that the randomization might be implemented in three distinct ways, by conditioning on different parts of the observed data:
  1. R1.
    Condition on $\rho(X)_{1:n}$ and $\tilde{Y}_{1:n}$, with $\tilde{Y}_{i}=\tilde{\tau}^{-1}_{i}Y_{i}$, and where if necessary $\tilde{\tau}_{i}$ is sampled from the inversion kernel and fixed. Set $\mathsf{\tilde{Y}}_{i}^{(b)}=\tilde{Y}_{i}$ and sample $\mathsf{\Gamma}_{i}^{(b)}\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$.
  2. R2.
    Condition on $\rho(X)_{1:n}$ and $\tilde{\tau}_{1:n}$, where if necessary the latter are sampled from the inversion kernel and fixed. Set $\mathsf{\Gamma}_{i}^{(b)}=\tilde{\tau}_{i}$ and sample $\mathsf{\tilde{Y}}_{i}^{(b)}\sim P_{\tilde{Y}|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$.
  3. R3.
    Condition on $\rho(X)_{1:n}$. Sample both $\mathsf{\Gamma}_{i}^{(b)}\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$ and $\mathsf{\tilde{Y}}_{i}^{(b)}\sim P_{\tilde{Y}|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(X_{i}))$.
Each of the procedures yields a different test; we differentiate between them by indicating what statistic of $(X,Y)_{1:n}$ is conditioned on, denoted by $S_{1:n}\coloneq S(X,Y)_{1:n}$. For example, in R1, $S_{i}=(\rho(X_{i}),\tilde{Y}_{i})$. Carrying out the appropriate randomization for each $i=1,\dotsc,n$ and repeating $B$ times, the CRT with test statistic $T$ estimates a conditional $p$-value as
$p_{B}\coloneqq\frac{1+\sum_{b=1}^{B}\mathbbm{1}\left\{T\left((X,Y)_{1:n}^{(b)}\right)\geq T\left((X,Y)_{1:n}\right)\right\}}{B+1}\;.$ (15)
The decision to reject the null hypothesis is taken if $p_{B}$ is less than or equal to a specified significance $\alpha\in[0,1]$. The test statistic $T$ should be some signal of distributional equality, with small values constituting evidence in favor of the null hypothesis. For example, we implement our tests in the subsequent sections using kernel maximum mean discrepancy (MMD).
It is straightforward to show via Proposition 4.2 and its corollaries that the resulting CRT based on $B$ randomization sets is a level-$\alpha$ test. Moreover, as $B\to\infty$, the CRT produces a consistent conditional $p$-value, in the following sense. Under the null hypothesis that $P_{Y|X}$ is $\mathbf{G}$-equivariant, if $(X^{\prime},Y^{\prime})\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(X,Y)$, then for $P$-almost every sequence $s_{1:n}$,
$p_{B}\;\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{\to}}\;\mathbb{P}\left(T\left((X^{\prime},Y^{\prime})_{1:n}\right))\geq T\left((X,Y)_{1:n}\right)\mid S^{\prime}_{1:n}=S_{1:n}=s_{1:n}\right)\;.$ (16)
Proposition 5.1 formalizes the result generally in terms of a test based on Proposition 4.2; the same result holds for a test based on Corollary 4.3. The proof is straightforward and provided in Section B.3 for completeness. As with the test of distributional invariance in Section 2.2, conditioning on $S_{1:n}$ means that $S_{1:n}$ can be used as inputs to the test statistic.
Proposition 5.1.
Assume that $P_{Y|X}$ is $\mathbf{G}$-equivariant. For data $(X,Y)_{1:n}$, generate $B$ randomized versions $(X,Y)^{(b)}_{1:n}$ using any of the randomization procedures R1R3. Assume that $\mathbb{P}(T((X,Y)_{1:n}^{(b)})=T((X,Y)_{1:n}^{(b^{\prime})}))=0$ for $b\neq b^{\prime}$. For arbitrary fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$, the statistic $p_{B}$ defined in (15) satisfies
[…]
Moreover, the conditional strong consistency of $p_{B}$, as in (16), holds.
5.1 Approximate conditional randomization
In Appendix C, we describe certain settings that allow for exact sampling from $P_{\Gamma|[X]}$. One particular case of interest described there is when $\mathbf{G}$ acts transitively on $\mathbf{X}$, in which case sampling from $P_{\Gamma|[X]}$ amounts to permuting the $\tilde{\tau}_{i}$’s. If $\mathbf{G}$ induces a finite number of orbits in $\mathbf{X}$, then an orbit-stratified permutation can be used.
In general, we expect that sampling from approximations of $P_{\Gamma|[X]}$ and/or $P_{\tilde{Y}|[X]}$ will be the more common approach in practice. To that end, let $(X,Y)_{i,m}^{(b)}$ be a randomization of $(X,Y)_{i}$ generated as in Proposition 5.1, except that $\mathsf{\Gamma}_{i}^{(b)}$ and/or $\mathsf{\tilde{Y}}_{i}^{(b)}$ is sampled from an approximation to the appropriate conditional distribution. For example, this might be a sample from a kernel density estimate constructed from an independent dataset of $m$ observations. Under the null hypothesis, if as $m\to\infty$, $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)$ conditioned on $\rho(X_{i})$, then the approximate CRT will be asymptotically level-$\alpha$.
Proposition 5.2.
Assume $(X,Y)\sim P_{X,Y}$, with $P_{Y|X}$ $\mathbf{G}$-equivariant. Let $T$ be a continuous $\mathbb{R}$-valued function, such that the distribution of $T((X,Y)_{1:n})$ is atomless. Suppose that $(X,Y)_{i,m}^{(b)}$ are sampled conditionally on $S(X,Y)_{i,m}^{(b)}=S(X,Y)_{i}$, and used as the randomization sets to generate CRT statistics as in (15), denoted $p_{B,m}$. If, for each $i\leq n$, conditioned on $S(X,Y)_{i}$, $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)_{i}$ as $m\to\infty$, and $\mathbb{P}(T((X,Y)_{1:n}^{(b)})=T((X,Y)_{1:n}^{(b^{\prime})}))=0$ for $b\neq b^{\prime}$, then $p_{B,m}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}p_{B}$. Therefore, for fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$,
[…]
The proof can be found in Section B.3. We note that the sequence of approximations do not necessarily require an independent dataset to estimate the conditional distributions, but can depend on $(X,Y)_{1:n}$ through $S_{1:n}$. In that case, the limit in Proposition 5.2 should be taken as $n\to\infty$.
Any asymptotically correct method for approximating the required conditional distributions can be used for the conditional randomization procedure to yield an asymptotically level-$\alpha$ test. Here, we describe a kernel estimation approach for sampling from approximations of the conditional distributions required for R1, R2 and R3.
Let $M(X)$ be any maximal invariant to condition on, such as $\rho(X)$, and let $k_{h_{m}}\colon\mathbf{M}\times\mathbf{M}\to\mathbb{R}_{+}$ be a kernel function defined on $\mathbf{M}$. Assume that $k_{h_{m}}$ is bounded above by $\kappa<\infty$, with $k_{h_{m}}(z,z)=\kappa$ for all $z\in\mathbf{M}$. Using a (possibly independent) dataset of size $m$ for estimation, assume that the bandwidth parameter $h_{m}$ decreases to 00 as $m\to\infty$, slowly enough to ensure asymptotically consistent density estimation. To generate a randomization conditioned on $M(X)$, we first assign to each index $i\in\{1,\dotsc,m\}$ a probability proportional to $k_{h_{m}}(M(X),M(X_{i}))$. To sample from an approximation of $P_{\tilde{Y}|[X]}$ for R2 and R3, we then sample an index $\ell$ from $\{1,\dotsc,m\}$ with those probabilities, and then take $\mathsf{\tilde{Y}}_{i}=\tilde{Y}_{\ell}$. Similarly, to sample from an approximation of $P_{\Gamma|[X]}$ for R1 and R3, we sample an index $j$ from $\{1,\dotsc,m\}$ (independent of $\ell$) with those probabilities, and then take $\mathsf{\Gamma}_{i}=\tilde{\tau}$, where $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X_{j})$. It can be seen that as $m\to\infty$, the distributions from which this procedure samples approach the target conditional distributions. The resulting conditional randomization procedure (without an independent dataset) used in a test based on Properties E1, E2 and E3 is summarized in Algorithm 2. For a test based on Properties E4, E5 and E6, only the second component of the output pair is retained.
Algorithm 2 Approximate conditional randomization procedure
1:inputs $(X,Y)$, $(X,Y)_{1:n}$
2:compute $p_{i}\coloneq k_{h_{m}}(M(X),M(X_{i}))$ for $i\in\{1,\dotsc,n\}$
3:independently sample $j,\ell\in\{1,\dotsc,n\}$ with probability proportional to $p_{1},\dotsc,p_{n}$
4:if randomize $\Gamma$ then
5:     sample $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X_{j})$
6:else
7:     sample $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X)$
8:end if
9:set $\mathsf{\Gamma}\coloneq\tilde{\tau}$
10:if randomize $\tilde{Y}$ then
[…]
12:else
[…]
14:end if
15:return $(\mathsf{\Gamma}\rho(X),\mathsf{\Gamma}\mathsf{\tilde{Y}})$
In our experiments, we estimate the conditional distributions using the same data used for testing. This procedure remains valid because we base the tests on Property E4, using only $\tilde{Y}_{1:n}$ and $Y_{1:n}$; the part of the observed data used for estimating conditional distributions, $\tilde{\tau}_{1:n}$, is conditionally independent of those variables under the null hypothesis. When the maximal invariant is continuous, we find that is it sufficient to use a Gaussian kernel with a diagonal covariance matrix given by Silverman’s rule-of-thumb (Silverman, 1986, Ch. 4).
5.2 Baseline: two-sample permutation tests
As a simple alternative to the conditional randomization tests described in the previous sections, consider generating $B=1$ randomized sample, then comparing it to the observed data using a two-sample test for equality in distribution. For example, a permutation test for equality in distribution provides a nonparametric baseline. Although these are valid level-$\alpha$ tests, they do not take full advantage of the structure induced by the group symmetry and we do not expect them to be as powerful against alternatives as the CRT. We implement such tests for baseline comparison in our experiments.
6 Kernel hypothesis tests
In the experiments reported in Section 7, we test for conditional symmetry using a kernel-based test statistic on observations denoted by $Z_{1:n}$. For tests based on Properties E1, E2 and E3, $Z_{i}=(X,Y)_{i}$, with a kernel defined on the product space $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$. For tests based on Properties E4, E5 and E6, $Z_{i}=Y_{i}$, with a kernel defined on $\mathbf{Z}=\mathbf{Y}$. We provide a brief overview of kernel-based hypothesis testing, and other related ideas in this section. See Christmann and Steinwart (2008) for a thorough treatment of the theory of kernels, and Muandet et al. (2017) for a review of their uses in statistics and machine learning.
Let $k_{\sigma}\colon\mathbf{Z}\times\mathbf{Z}\rightarrow\mathbb{R}$ be a positive semidefinite kernel with parameter $\sigma$, corresponding to the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$. If the kernel is measurable and bounded, then the kernel mean embedding (KME) of a probabililty measure $P$ on $\mathbf{Z}$ is the unique element $\mu_{P}({\,\vbox{\hbox{\tiny$\bullet$}}\,})\coloneqq\int_{\mathbf{Z}}k_{\sigma}(z,{\,\vbox{\hbox{\tiny$\bullet$}}\,})P(dz)\in\mathcal{H}$ such that $\mathbb{E}_{P}[f(Z)]=\langle f,\mu_{P}\rangle_{\mathcal{H}}$ for all $f\in\mathcal{H}$ (Gretton et al., 2012). It follows that $\langle\mu_{P_{1}},\mu_{P_{2}}\rangle=\int_{\mathbf{Z}\times\mathbf{Z}}k_{\sigma}(z,z^{\prime})P_{1}(dz)P_{2}(dz^{\prime})$. A kernel is said to be characteristic if the map $P\mapsto\mu_{P}$ from $\mathcal{P}(\mathbf{Z})$ into $\mathcal{H}$ is injective, meaning that the embedding for each probability measure is unique (Sriperumbudur et al., 2010). Kernel-based hypothesis tests compare distributions through their KMEs (Gretton et al., 2007). The test statistic in kernel-based tests is an estimator of the squared maximum mean discrepancy (MMD),
$\mathrm{MMD}_{\sigma}^{2}(P_{1},P_{2})\coloneqq\left\|\mu_{P_{1}}-\mu_{P_{2}}\right\|_{\mathcal{H}}^{2}$
If the kernel $k_{\sigma}$ is characteristic, $\mathrm{MMD}_{\sigma}^{2}$ is a metric on $\mathcal{P}(\mathbf{Z})$ (Gretton et al., 2012); hence $\mathrm{MMD}_{\sigma}(P,Q)=0$ if and only if $P=Q$. If a non-characteristic kernel is used instead, $\mathrm{MMD}_{\sigma}^{2}$ would be unable to separate some distinct elements of $\mathcal{P}(\mathbf{Z})$, which would not affect the level of the test, but may reduce the power under some alternatives.
Algorithm 3 Test for conditional symmetry using MMD-based statistics
1:inputs: $(X,Y)_{1:n}$, $B$, $\alpha$, reuse option, statistic $T\in\{\widehat{\mathrm{MMD}}^{2}_{\sigma},\widehat{\mathrm{FUSE}}_{\omega},\widehat{\textrm{SK}}\}$, $\{\sigma_{1},\dotsc,\sigma_{L}\}$
2:resample $Z^{(0)}_{1:n}$ using Algorithm 2
3:compute $T^{(0)}\coloneq T(Z_{1:n},Z_{1:n}^{(0)})$
4:if reuse then
5:     resample $Z^{\prime}_{1:n}$ using Algorithm 2
6:end if
7:for $b$ in $1\dotsc,B$ do
8:     resample $Z^{(b)}_{1:n}$ using Algorithm 2
9:     if !reuse then
10:         resample $Z^{\prime}_{1:n}$ using Algorithm 2
11:     end if
12:     compute $T^{(b)}\coloneq T(Z^{(b)}_{1:n},Z^{\prime}_{1:n})$
13:end for
14:compute $p_{B}$ as
$p_{B}\coloneqq\frac{1+\sum_{b=1}^{B}\mathbbm{1}\{T^{(b)}\geq T^{(0)}\}}{1+B}$
15:reject $H_{0}$ if $p_{B}\leq\alpha$
The MMD can be estimated from independent samples $Z_{1:n}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P_{1}$ and $Z_{1:n}^{\prime}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P_{2}$ via the unbiased U-statistic
$\widehat{\mathrm{MMD}}_{\sigma}^{2}(Z_{1:{n}},Z_{1:{n}}^{\prime})=\frac{1}{n(n-1)}\sum_{i\neq j}k_{\sigma}(Z_{i},Z_{j})+k_{\sigma}(Z_{i}^{\prime},Z_{j}^{\prime})-\frac{2}{n^{2}}\sum_{i,j}k_{\sigma}(Z_{i},Z_{j}^{\prime})\;,$
or the biased V-statistic
$\widehat{\mathrm{MMD}}_{v,\sigma}^{2}(Z_{1:{n}},Z_{1:{n}}^{\prime})=\frac{1}{n^{2}}\sum_{i,j}k_{\sigma}(Z_{i},Z_{j})+k_{\sigma}(Z_{i}^{\prime},Z_{j}^{\prime})-2k_{\sigma}(Z_{i},Z_{j}^{\prime})\;.$
Our CRT implementation, shown in Algorithm 3, is able to use either of these statistics, or the adaptive MMD-based statistics described in the next section. The basic idea is to simulate conditionally independent randomizations $Z_{1:n}^{(b)}$ and ${Z^{\prime}}_{1:n}^{(b)}$, then use them to estimate the MMD. This produces a null distribution with which to compare the estimated MMD between $Z_{1:n}$ and yet another independent randomization ${Z}^{(0)}_{1:n}$. Optionally, a single set of comparison samples $Z^{\prime}_{1:n}$ can be reused (option reuse in Algorithm 3) instead of independent comparison samples ${Z}_{1:n}^{(b)}$, $b=0,\dotsc,B$. The level of the test in each case is $\alpha$, though we expect different power properties; we study both versions in the experiments in Section 7.
In the theoretical analysis of the MMD-based tests, not all of the usual statistical properties of MMD-based tests hold. Whereas in the standard setting independent samples are compared, here the samples are dependent through conditioning on $S_{1:n}$ for the randomization. The two samples are, however, conditionally independent, which allows us to build on established properties of MMD estimators when independent samples are used. For ease of theoretical analysis of the MMD-based test, we assume that independent comparison sets (reuse = false) and the V-statistic are used. We use the U-statistic in our experiments, but we do not expect the two to behave very differently.
Conditioned on $S_{1:n}$, the expected value of the V-statistic is
$\Delta_{n}^{2}\coloneq\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\;\bigg{|}\;S_{1:n}\right]$ (18)
$=\frac{1}{n^{2}}\sum_{i,j}\mathbb{E}\left[k_{\sigma}(Z_{i},Z_{j})+k_{\sigma}(Z^{(0)}_{i},Z^{(0)}_{j})-2k_{\sigma}(Z_{i},Z^{(0)}_{j})\mid S_{i},S_{j}\right]\;.$
If $\Delta_{n}$ is sufficiently large, we expect the test to have power against non-equivariant alternatives. To state the formal result, let $\nu^{2}=\sup_{z}k_{\sigma}(z,z)<\infty$, and define for $\eta>0$,
$R_{n}(\eta)\coloneq\frac{9}{2n}\nu^{2}+\frac{16\nu^{2}\sqrt{2}}{\sqrt{n^{3}}}+6\nu\sqrt{\frac{2}{n}}+\frac{3\nu}{2}\sqrt{\eta}\sqrt{\frac{2}{n}+32\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta\nu}{n}\;.$ (19)
Because we analyze the conditional (on $S_{1:n}$) test, probabilistic properties of the test can be said to hold conditionally (on $S_{1:n}$), almost surely (with respect to $S_{1:n}$). Analysis of bootstrap procedures rely on the same ideas (e.g., van der Vaart, 1998, Ch. 23).
Theorem 6.1.
Let $T$ be $\widehat{\mathrm{MMD}}_{v,\sigma}^{2}$, with kernel $k_{\sigma}$ and corresponding RKHS $\mathcal{H}_{k_{\sigma}}$, using exact randomization in the CRT. Assume that $\Delta_{n}\geq 2\nu\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-e^{-\eta}$ over observations $Z_{1:n}$ and randomization ${Z}^{(0)}_{1:n}$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\Delta_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;,$ (20)
where $F_{B,p(\Delta_{n})}$ is the cumulative distribution function for the binomial distribution of $B$ trials with success probability
$p_{\eta}(\Delta_{n})=\exp\left(-\frac{3}{2}\Delta_{n}+R_{n}(\eta)\right)\;.$ (21)
The proof can be found in Section B.4. Observe that for fixed $\alpha$, $B$, and $n$, the lower bound in (20) is increasing in $\Delta_{n}$. Because multiple approximations were used for the sake of deriving explicit bounds, tighter lower bounds seem possible for (21) and in Corollary 6.2 below, but we do not pursue them here.
Using the Chernoff bound on the upper tail of deviations of a binomial random variable, we can obtain an explicit lower bound on values of $\Delta_{n}$ that correspond to alternatives for which the test (with high probability) rejects the null with probability at least $\beta\in(0,1)$. For convenience, let $B_{\alpha}\coloneq\lfloor\alpha(B+1)-1\rfloor$.
Corollary 6.2.
Assume the conditions of Theorem 6.1 and that $B_{\alpha}\geq Bp_{\eta}(\Delta_{n})$. Fix any $\beta\in(0,1)$. For each $\eta>0$, if
$\Delta_{n}\geq\frac{2}{3}\left(R_{n}(\eta)-\ln\frac{B_{\alpha}}{B}-\ln\left(1-\sqrt{\frac{-2\ln(1-\beta)}{B_{\alpha}}}\right)\right)\;,$ (22)
then for almost every sequence $(S_{i})_{i\geq 1}$, conditionally on $S_{1:n}$, we have that $\mathbb{P}(p_{B}\leq\alpha|S_{1:n})\geq\beta$ with probability over $(Z_{1:n},{Z}^{(0)}_{1:n})$ at least $1-e^{-\eta}$.
In the limit of $B,n\to\infty$, we see that if $\text{MMD}(P,P^{(0)})\geq-\frac{2}{3}\ln\alpha$, then the test will reject with probability $\beta$ for any $\beta\in(0,1)$.
6.1 Adaptive kernel test statistics
It is known that the power of kernel-based tests are highly dependent on the chosen bandwidth parameter $\sigma$ (Ramdas et al., 2015). Several solutions have been proposed, in the form of data-based heuristics (Gretton et al., 2012; Garreau, Jitkrittum and Kanagawa, 2017) and, more recently, adaptive methods (Schrab et al., 2023; Biggs, Schrab and Gretton, 2024; Cárcamo, Cuevas and Rodríguez, 2024). In our experiments, when using a kernel with a parameter (e.g., the Gaussian kernel), we use two test statistics based on solutions proposed in the adaptive kernel literature. The first adapts the unnormalized FUSE statistic
$\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},Z_{1:n}^{\prime})=\frac{1}{\omega}\log\left(\frac{1}{L}\sum_{\ell=1}^{L}\exp\left(\omega\widehat{\text{MMD}}^{2}_{\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)\right)\;,$
where in practice $\omega=\sqrt{n(n-1)}$ following Biggs, Schrab and Gretton (2024). The statistic can be interpreted as taking a softmax across estimated MMD2 values computed using different kernels $k_{\sigma_{\ell}}$. The bandwidth values $\{\sigma_{1},\dotsc,\sigma_{L}\}$ in the original work (Biggs, Schrab and Gretton, 2024) are quantiles of the pairwise Euclidean distances in the data. In the context of a permutation test, the empirical measure of the data is a maximal invariant, so any function of it (such as quantiles) can be used in the context of a test for permutation-invariance, as in Theorem 2.2. Our adaptation, which satisfies the requirements of Proposition 5.1, is obtained by replacing the quantiles of the observation with quantiles of variables contained in the conditioning set $S_{1:n}$, such as the pairwise Euclidean distances in $\tilde{Y}_{1},\dotsc,\tilde{Y}_{n}$.
The second test statistic we use is an adaptation of the supremum kernel distance (SK, Cárcamo, Cuevas and Rodríguez 2024) given by
$\widehat{\mathrm{SK}}(Z_{1:n},Z_{1:n}^{\prime})=\sup_{\ell\in\{1,\dotsc,L\}}\widehat{\text{MMD}}^{2}_{\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\;.$
Tests using this procedure remain valid due to the conditioning argument in Proposition 5.1.
Analogous versions of the power lower bound in Theorem 6.1 hold for the FUSE and SK statistics. We state a combined version here; separate proofs are given in Proposition B.7 and Proposition B.8, respectively, in Section B.5. For convenience, let $R_{n}^{\vee}(\eta)$ denote the maximum over $\ell\in[L]$ of $R_{n,\ell}(\eta)$, which is obtained by substituting $\nu_{\ell}$ for $\nu$ into (19). Define
$\bar{\Delta}_{n}\coloneq\frac{1}{L}\sum_{\ell=1}^{L}\sqrt{\mathbb{E}\left[\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z^{\prime(0)}_{1:n})\mid S_{1:n}\right]}\;,$
and
$\Delta_{n}^{\vee}=\sup_{\ell\in[L]}\Delta_{n,\ell}\coloneq\sup_{\ell\in[L]}\sqrt{\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\mid S_{1:n}\right]}\;.$
Proposition 6.3.
Let $T$ be the FUSE (resp. SK) statistic using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. Assume that for each $\ell\in\{1,\dotsc,L\}$, $\Delta_{n,\ell}\geq 2\nu_{\ell}\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-Le^{-\eta}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\widetilde{\Delta}_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
where $F_{B,p_{\eta}(\Delta_{n}^{\vee})}$ is the distribution function for the binomial distribution of $B$ trials with success probability
$p_{\eta}(\widetilde{\Delta}_{n})=\min\left\{1,L\exp\left(-\frac{3}{2}\widetilde{\Delta}_{n}+R_{n}^{\vee}(\eta)\right)\right\}\;,$
and $\widetilde{\Delta}_{n}=\bar{\Delta}_{n}$ for FUSE, and $\widetilde{\Delta}_{n}=\Delta^{\vee}_{n}$ for SK.
Using this, a lower bound for $\widetilde{\Delta}_{n}$ that guarantees power $\beta$ with high probability, analogous to the one in Corollary 6.2, can be obtained.
7 Experiments
We evaluate our proposed tests on simple synthetic examples and on realistic simulated data from particle physics. In all of our experiments, we sample $n$ data points from a distribution or dataset and perform a test for a conditional symmetry with respect to some specified group. We repeat this procedure over $N=1000$ simulations for each test setting and report the proportion of simulations in that the test rejected, which is an estimate of the test size (resp. power) when the distribution has (resp. does not have) the specified symmetry. With $N=1000$ simulations, estimates are precise up to approximately ±$\pm 0.016$. We set the test level to be $\alpha=0.05$ in all of our experiments. Where applicable, we use $B=100$ randomizations for estimating the $p$-value in randomization tests, except when specified otherwise.
We default to aggregating Gaussian and Laplace kernels via the FUSE and SK test statistics for data that are continuous. To compute the bandwidths $\{\sigma_{1},\dotsc,\sigma_{L}\}$, we follow a procedure similar to the one described by Biggs, Schrab and Gretton (2024). We first compute the distances $\{\|\tilde{Y}_{i}-\tilde{Y}_{j}^{\prime}\|_{2}\}_{i,j}$ and $\{\|\tilde{Y}_{i}-\tilde{Y}_{j}^{\prime}\|_{1}\}_{i,j}$ for the Gaussian and Laplace kernels, respectively. We then compute half of the 5% quantile and two times the 95% quantile for each set, and take the $L=10$ bandwidth values to be the discrete uniform interpolation of those two values. All estimates of MMD use the unbiased U-statistic.
All experiments were implemented in Julia (version 1.6.1) and run on a high-performance computing allocation with 4 CPU cores and 16 GB of RAM. Additional experiments and experimental details can be found in Appendix D. Associated computer code can be found on GitHub. [^3] https://github.com/chiukenny/Tests-for-Conditional-Symmetry
7.1 Experimental configurations
In our experiments involving tests for conditional symmetry, we consider up to eight tests, each of which corresponds to a particular configuration of the following parameters:
  1. 1.
    Permutation test (BS; Section 5.2) versus conditional randomization test (RN; Section 5).
  2. 2.
    Reusing a randomized comparison set (reuse = true in Algorithm 3) versus using independent comparison sets (reuse = false).
  3. 3.
    FUSE test statistic (FS) versus supremum kernel test statistic (SK).
For example, the permutation test that uses the FUSE test statistic is denoted BS-FS. We also consider a similar baseline when conducting a test for joint invariance as a test for conditional symmetry, where the samples used to estimate the $p$-value are generated via permutations rather than group actions from $\lambda$. In this setting, we evaluate two tests: the permutation test described here (BS) and the group-based randomization test (RN; Section 2.2).
Figure 6: Rejection rates for tests that reuse a comparison set in the changing covariance experiment for varying values of the data dimension $d$, $Y|X$ covariance matrix off-diagonal values $p$, and sample size $n$. The conditional distribution is $\mathrm{SO}(d)$-equivariant only for $p=0$.
7.2 Synthetic experiments
We start with several synthetic experiments involving simple distributions for the marginal of $X$ and the conditional $Y|X$. These experiments will be used to compare and demonstrate the properties of the different tests for conditional symmetry. One of these experiments is discussed in this section; the rest can be found in Appendix D.
Figure 7: Rejection rates for tests that reuse a comparison set in the changing covariance experiment, where the non-$\mathrm{SO}(2)$-equivariance comes from the covariance $p$ between the first two dimensions, for varying values of the data dimension $d$ and sample size $n$.
Consider random variables $X$ and $Y$ in $\mathbb{R}^{d}$. Let $X$ have a marginal distribution given by $P_{X}=\mathsf{N}(0,\Sigma_{X,d})$, where $\Sigma_{X,d}$ is the $d\times d$ matrix with $1$’s on the diagonal and $0.75$ on the off-diagonals. Let $Y|X$ have the conditional distribution $\mathsf{N}(X,\Sigma_{Y,d}(p))$, where $\Sigma_{Y,d}(p)$ is the $d\times d$ matrix with $1$’s on the diagonals and $p$ on the off-diagonals. In words, the distribution of $Y|X$ is a multivariate normal centered on $X$. For $p=0$, the distribution is spherical; as $p$ moves towards $1$, the distribution becomes more elliptical and eventually line-like. Under the action of $\mathrm{SO}(d)$, the conditional distribution is $\mathrm{SO}(d)$-equivariant if and only if $p=0$.
To test for $\mathrm{SO}(d)$-equivariance on data generated from these distributions, we condition on the maximal invariant $M(X)=\|X\|$. Here, we only show results for the tests that reuse a comparison set; results for tests that use independent comparison sets in this experiment can be found in Section D.2. Figure 6 shows the estimated test rejection rates for various values of $p$ and $d$ as the sample size $n$ increases from $25$ to $500$. The randomization tests appear to be more powerful than the baseline tests in general. We also observe that increasing the dimensions leads to an increase in power, though this appears to be an artifact of the experimental setup. Consider a similar setup but where the only non-zero off-diagonal entry of $\Sigma_{Y,d}$ is between the first and second dimensions. In other words, the non-equivariance of the distribution is isolated to the first two dimensions. Figure 7 shows the rejection rates for this setting. Here, the power decreases with dimension, which suggests that the non-equivariance signal-to-noise ratio decreases with additional “equivariant dimensions”. Histograms of the $p$-values in these experiments can be found in Section D.2.
Additional synthetic experiments are found in Appendix D: under a similar setting, Section D.3 investigates the effect of the number of randomizations $B$ and Section D.4 compares sampling from an approximate conditional distribution to sampling from the true conditional; Section D.5 investigates non-equivariance in the mean as opposed to in the covariance; Section D.6 investigates tests for permutation-equivariance in another synthetic setting.
7.3 Large Hadron Collider dijet events
Table 1: Rejection rates (over $1000$ repetitions) for tests for $\mathrm{SO}(2)$-equivariance on LHC dijet data samples of size $n=100$.
Test
Comparison set BS-FS BS-SK RN-FS RN-SK
Original data Reused $0.000$ $0.002$ $0.000$ $0.000$
Independent $0.002$ $0.003$ $0.005$ $0.004$
Shuffled data Reused $0.818$ $0.836$ $0.922$ $0.905$
Independent $0.838$ $0.815$ $0.901$ $0.889$
Our first application demonstrates how our tests could be used to verify symmetry presumed by theory and to detect possible instances of symmetry breaking for scientific discovery. The Large Hadron Collider (LHC) Olympics 2020 dataset (Kasieczka et al., 2021; Kasieczka, Nachman and Shih, 2019) consists of 1.1 million simulated dijet events generated by PYTHIA (Bierlich et al., 2022), a widely-used Monte Carlo simulator for high-energy physics processes. A dijet is two jets of constituent particles produced from the decay of high-energy quarks following a particle collision. The transverse momentum, polar angle $\phi$ and pseudorapidity $\eta$ for up to 200 constituents were recorded for each jet; the Cartesian momentum in the transverse plane for each constituent can be computed from these values as the pair $(p_{1}=p_{\text{T}}\cos(\phi),p_{2}=p_{\text{T}}\sin(\phi))$. The leading constituents of a jet are the particles with the largest transverse momenta in any direction.
By conservation of angular momentum, the Cartesian momenta of one leading constituent, $Y=(p_{1,Y},p_{2,Y})$, should be equivariant with respect to $\mathrm{SO}(2)$ rotations of the other, $X=(p_{1,X},p_{2,X})$. We conduct tests for $\mathrm{SO}(2)$-equivariance on random samples of size $n=100$ drawn from the full data. The estimated rejection rates are shown in Table 1. Regardless of whether a randomized comparison set is reused or not, we find that all tests achieve a rejection rate of no more than $0.005$, suggesting the presence of $\mathrm{SO}(2)$-equivariance, which is consistent with the theory. As a check for power, we also consider tests for $\mathrm{SO}(2)$-equivariance on samples of the same size where the second leading constituents were permuted between observations, breaking the dependence and thus the conditional equivariance. The tests all reject with a power greater than $0.8$, which suggests the absence of symmetry in this setting.
Figure 18 in Section D.7 visualizes the transverse momentum of the leading constituents. It can be observed that the momenta $X$ of the first leading constituents appear to be distributed on a ring centered at the origin, suggesting that the distribution of $X$ may be marginally invariant with respect to the action of $\mathrm{SO}(2)$. As per Property E7, equivariance could then be tested via a test for joint invariance under the action of the group of paired rotations $\mathbf{G}_{0}=\{(g_{1},g_{2})\in\mathrm{SO}(2)\times\mathrm{SO}(2):g_{1}=g_{2}\}$. We conduct a test for invariance with respect to this subgroup on samples of size $n=25$. Furthermore, as the presence of unexpected symmetry would also be of scientific interest, we also test for invariance under the action of the full $\mathbf{G}_{1}=\mathrm{SO}(2)\times\mathrm{SO}(2)$ group and of $\mathbf{G}_{2}=\mathrm{SO}(4)$. The results are shown in Table 2. We see that our tests detect $\mathbf{G}_{0}$-invariance and reject $\mathbf{G}_{1}$- and $\mathbf{G}_{2}$-invariance, which is again consistent with the Standard Model.
Table 2: Rejection rates (over $1000$ repetitions) for tests for joint invariance on LHC dijet data samples of size $n=25$.
Test
Group to test Comparison set BS-FS BS-SK RN-FS RN-SK
[…]
Independent $0.044$ $0.048$ $0.059$ $0.040$
[…]
Independent $0.425$ $0.838$ $0.946$ $0.970$
$\mathrm{SO}(4)$ Reused $0.615$ $0.915$ $0.983$ $0.976$
Independent $0.611$ $0.917$ $0.982$ $0.977$
7.4 Lorentz symmetry of quark decays
We consider a second particle physics application based on the Top Quark Tagging Reference dataset (Kasieczka et al., 2019), which also consists of jet events simulated by PYTHIA. The original dataset was constructed for the task of classifying jet events as having decayed from a top quark or not and consists of a training, validation, and test set. The four-momenta $x=(E,p_{1},p_{2},p_{3})$ of up to 200 jet constituents ordered by transverse momentum (i.e., highest magnitude momentum in the plane perpendicular to the particle beam line) are recorded for each event. We only use the test set, which contains 404,000 simulated events, and ignore the binary label that indicates whether the jet decayed from a top quark or not. According to the Standard Model, the four-momenta of the constituents in a jet should be equivariant with respect to the Lorentz group $\mathrm{SO}^{+}(1,3)$, which was described in Examples 4 and 6, and preserves the quadratic form $Q(x)=E^{2}-p_{1}^{2}-p_{2}^{2}-p_{3}^{2}$.
We take $X$ to be the four-momentum of the leading constituent in each jet, and $Y$ to be that of the second leading constituent (as in Yang et al., 2023). We conduct tests for $\mathrm{SO}^{+}(1,3)$-equivariance of $P_{Y\mid X}$ based on samples of size $n=100$ using the maximal invariant $Q(x)$. The estimated rejection rates are shown in Table 3. The tests all reject $\mathrm{SO}^{+}(1,3)$-equivariance at a rate of less than $0.01$, indicating symmetry consistent with the Standard Model. Like in the LHC dijet experiment, we also conduct tests on samples where the second leading constituents are permuted across the data. In this setting, the conditional randomization tests are able to detect the breaking of $\mathrm{SO}^{+}(1,3)$-equivariance with reasonable power. The baseline permutation tests are not as powerful in this setting.
Table 3: Rejection rates (over $1000$ repetitions) for tests for $\mathrm{SO}^{+}(1,3)$-equivariance on quark decay data samples of size $n=100$.
Test
Comparison set BS-FS BS-SK RN-FS RN-SK
Original data Reused $0.000$ $0.001$ $0.002$ $0.004$
Independent $0.000$ $0.001$ $0.006$ $0.009$
Shuffled data Reused $0.281$ $0.195$ $0.780$ $0.752$
Independent $0.281$ $0.195$ $0.686$ $0.650$
7.5 Training an invariant MNIST classifier
As our final application, we illustrate how our tests for symmetry of a conditional distribution $P_{Y|X}$ can be adapted as tests for symmetry of a function $f:\mathbf{X}\to\mathbf{Y}$ by taking $P_{Y|X}=\delta_{f(X)}$. In light of this, our tests could be used as, for example, validation tools for assessing whether a model has learned symmetry after being trained on data. These test-based tools would only require data in the form of input and predicted output pairs. To demonstrate this utility, we consider a setting where a MNIST classifier is trained to be invariant with respect to rotations of MNIST images via data augmentation (Chen, Dobriban and Lee, 2020; Huang, Orbanz and Austern, 2022). Data augmentation aims to desensitize the model to certain transformations by randomly transforming the training data. Therefore, unlike methods that hardcode symmetry as inductive biases, it may not always be obvious whether a model has learned symmetry until after it is applied in practice.
The MNIST dataset is a collection of $28\times 28$ pixel images of handwritten digits. The training set consists of $20000$ images; the test set consists of $45000$ augmented images, from which we uniformly sample to use in our tests. We train two identical LeNet convolutional neural networks for predicting the digit in the image, where one model is trained on images augmented with rotations and the other without. Before and during training, we pause the training process after every three epochs to test the invariance properties of each model. To generate the data used in each test, a sample of $n=100$ augmented images is first sampled from the test set and then fed into the model under examination. The data $Y$ are then taken as the digit probabilities predicted by the model, and so each observation is a vector on a simplex of dimension $d=10$ (or $d=9$ in the setting where the $9$ images is removed from the data). In our tests, we take the maximal invariant to be the unaugmented image to avoid complications with orienting images.
As augmentation ensures that the marginal distribution of the images will be invariant, we test for conditional invariance using tests for invariance via Property E8. For the test statistic, we use the information diffusion kernel for simplex data (Lebanon and Lafferty, 2002), given by
$k(y,y^{\prime})\approx(4\pi\sigma)^{-\frac{d+1}{2}}\exp\left(-\frac{1}{\sigma}\arccos^{2}\left(\textstyle\sum_{i=1}^{d}\sqrt{y_{i}y_{i}^{\prime}}\right)\right)\;,$
where $\sigma$ is the kernel parameter. To perform randomization in these tests, we apply random rotations to the images and then predict their label probabilities using the model trained until that point in the training process. We repeat this procedure for $N=1000$ instances at each testing point to estimate the rejection rate of the test, which can be interpreted as a measure of how invariant the model is at that point during training.
Figure 8: CRT rejection rates for image rotation-invariance of two MNIST models during training, where one is trained with data augmentation and the other without. The “R” prefix in the test label denotes reuse of a comparison set; the “I” prefix denotes independent comparison sets. The top row shows results for when the models are trained on the full MNIST data, and the bottom row shows results for when images with true label $9$ are removed from the data.
The estimated rejection rates for the randomization tests are shown in Figure 8 for two settings: one where the models are trained on the full MNIST data, and one where images with true label 9 are removed from the data. In both settings, the rejection rate for the model trained with augmentations is approximately $\alpha=0.05$ after the 30th epoch, whereas it is unclear if the rejection rate for the model trained without augmentations is converging to any value.
Here, we used the estimated test rejection rate as a measure of model invariance. It may be tempting to determine whether symmetry has been learned at a particular time point by whether a single test rejects or not, but conducting a suite of tests at multiple points during training would be subject to multiple testing considerations. A potential direction for future work is anytime-valid tests for conditional symmetry, which would allow for the continual evaluation of symmetry-learning based on a single test at each time point during training. Anytime-valid tests for marginal group invariance have recently appeared in the literature (Pérez-Ortiz et al., 2024; Lardy and Pérez-Ortiz, 2024).
Appendix A Background on groups
A group $\mathbf{G}$ is a set with a binary operation $\cdot$ that satisfies the associativity, identity, and inverse axioms. We denote the identity element by id. For notational convenience, we write $g_{1}g_{2}=g_{1}\cdot g_{2}$ for $g_{1},g_{2}\in\mathbf{G}$. A group $\mathbf{G}$ is said to be measurable if the group operations $g\mapsto g^{-1}$ and $(g_{1},g_{2})\mapsto g_{1}g_{2}$ are $\mathbf{S}_{\mathbf{G}}$-measurable, where $\mathbf{S}_{\mathbf{G}}$ is a $\sigma$-algebra of subsets of $\mathbf{G}$. A group action is a function $\Phi\colon\mathbf{G}\times\mathbf{X}\to\mathbf{X}$ that satisfies $\Phi(\textrm{id},x)=x$ and $\Phi(g_{1}g_{2},x)=\Phi(g_{1},\Phi(g_{2},x))$ for all $x\in\mathbf{X},g_{1},g_{2}\in\mathbf{G}$. A group $\mathbf{G}$ acts measurably on a set $\mathbf{X}$ if the group action $(g,x)\mapsto gx$ is measurable relative to $\mathbf{S}_{\mathbf{G}}\otimes\mathbf{S}_{\mathbf{X}}$ and $\mathbf{S}_{\mathbf{X}}$. Throughout, all actions are assumed to be continuous (and therefore measurable), and for convenience we simply say that $\mathbf{G}$ acts on $\mathbf{X}$, writing $gx=\Phi(g,x)$. This standard notation hides the particular action $\Phi$ associated with $gx$, and it is important not to lose track of that fact.
The stabilizer subgroup of $x\in\mathbf{X}$ is the subset of group actions that leave $x$ invariant, denoted $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$. The action is free (or exact) if $gx=x$ implies that $g=\textrm{id}$, in which case $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$. The $\mathbf{G}$-orbit of $x\in\mathbf{X}$ is the subset of all points in $\mathbf{X}$ reachable through a group action in $\mathbf{G}$, denoted $O(x)=\{gx:g\in\mathbf{G}\}$. Each orbit is an equivalence class of $\mathbf{X}$, where two points are equivalent if and only if they belong to the same orbit. The stabilizer subgroups of the elements of an orbit are all conjugate; that is, if $gx=x^{\prime}$ for $x\neq x^{\prime}$, then $g\mathbf{G}_{x}g^{-1}=\mathbf{G}_{x^{\prime}}$. If $\mathbf{X}$ has only one orbit, then the action is said to be transitive.
A.1 Proper group actions
In the analysis of probabilistic aspects of group actions, measurability issues can arise without regularity conditions. The key regularity condition that we assume is that the group action is proper: there exists a strictly positive measurable function $h\colon\mathbf{X}\to\mathbb{R}_{+}$ such that $\int_{\mathbf{G}}h(gx)\lambda(dg)<\infty$ for $x\in\mathbf{X}$ (Kallenberg, 2007). This definition of proper group action is a weaker, non-topological version of the definition commonly used in previous work in the statistics literature (e.g., Eaton, 1989; Wijsman, 1990; McCormack and Hoff, 2023), and only requires the existence of Haar measure. The common topological version defines $(g,x)\mapsto(gx,x)$ as a proper map, i.e., the inverse image of each compact set in $\mathbf{X}\times\mathbf{X}$ is a compact set in $\mathbf{G}\times\mathbf{X}$. That definition implies the one we use; see Kallenberg (2007) for details.
A sufficient condition for proper group action is that $\mathbf{G}$ is compact and acts continuously on $\mathbf{X}$. When $\mathbf{G}$ is non-compact, a group action can fail to be proper if $\mathbf{G}$ is “too large” for $\mathbf{X}$ in the sense that the stabilizer subgroups are non-compact. A class of non-compact group actions known to be proper are those of the isometry group of a non-compact Riemannian manifold. In this work, we rely on the assumption of proper group actions to guarantee the existence of measurable orbit selectors and inversion kernels. The proposition below compiles some known useful properties of these objects.
To state the proposition, let $\nu$ be any bounded measure on $(\mathbf{X},\mathbf{S}_{\mathbf{X}})$ and let $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$ be the completion of $\mathbf{S}_{\mathbf{X}}$ to include all subsets of $\nu$-null sets, and denote by $\bar{\nu}$ the extension of $\nu$ to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$ (see, e.g. Çinlar, 2011, Proposition 1.3.10). All statements of $\bar{\nu}$-measurability in the following proposition are with respect to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$, so that a set $A\subseteq\mathbf{X}$ is $\bar{\nu}$-measurable if $A\in\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$. Moreover, a function defined by a particular property is $\bar{\nu}$-measurable if it is measurable in the usual sense with respect to $\bar{\mathbf{S}}^{\nu}_{\mathbf{X}}$, and if the defining property holds with the possible exception of a $\bar{\nu}$-null set. Such a function would also be $\bar{\rho}$-measurable for any measure $\rho\ll\nu$.
Proposition A.1.
Let $\mathbf{G}$ be a lcscH group acting continuously and properly on $\mathbf{X}$, and $\nu$ any bounded measure on $\mathbf{X}$. Then the following hold:
  1. 1.
    The canonical projection $\pi\colon\mathbf{X}\to\mathbf{X}/\mathbf{G}$ is a measurable maximal invariant. Any measurable $\mathbf{G}$-invariant function $f\colon\mathbf{X}\to\mathbf{Y}$ can be written as $f=f^{*}\circ\pi$, for some measurable $f^{*}\colon\mathbf{X}/\mathbf{G}\to\mathbf{Y}$, and $f^{*}$ is bijective if and only if $f^{*}\circ\pi$ is a measurable maximal invariant.
  2. 2.
    There exists a $\bar{\nu}$-measurable orbit selector $\rho\colon\mathbf{X}\to[\mathbf{X}]$, which is a maximal invariant statistic, and it induces a $\bar{\nu}$-measurable cross-section $[\mathbf{X}]=\rho(\mathbf{X})$.
  3. 3.
    For a fixed $\bar{\nu}$-measurable orbit selector $\rho$, there exists a unique $\bar{\nu}$-measurable inversion probability kernel $P_{\tilde{\tau}|X}\colon\mathbf{X}\times\mathbf{S}_{\mathbf{G}}\to[0,1]$ with the following properties:
    1. (a)
      $P_{\tilde{\tau}|X}$ is $\mathbf{G}$-equivariant: For all $g\in\mathbf{G},x\in\mathbf{X},B\in\mathbf{S}_{\mathbf{G}}$, $P_{\tilde{\tau}|X}(B|gx)=P_{\tilde{\tau}|X}(g^{-1}B|x)$.
    2. (b)
      For each $x\in\mathbf{X}$, $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\rho(x))$ is normalized Haar measure on the stabilizer subgroup $\mathbf{G}_{\rho(x)}$.
    3. (c)
      For each $x\in\mathbf{X}$, if $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$, then $\tilde{\tau}\rho(x)=x$ with probability one.
    4. (d)
      If there is a $\bar{\nu}$-measurable representative inversion $\tau\colon\mathbf{X}\to\mathbf{G}$ associated with $\rho$ such that it satisfies $\tau(x)\rho(x)=x$ and $\tau(gx)=g\tau(x)$ for each $x\in\mathbf{X},g\in\mathbf{G}$, then $P_{\tilde{\tau}|X}^{\prime}(B|x)=P_{\tilde{\tau}|X}(\tau(x)^{-1}B|\rho(x))$ is an equivalent inversion kernel. In particular, this holds when the action of $\mathbf{G}$ on $\mathbf{X}$ is free, in which case $\mathbf{G}_{\rho(x)}=\{\textrm{id}\}$ and the inversion kernel is $\delta_{\tau(x)}$.
The measurability of the canonical projection is a result from functional analysis; see Eaton (1989, Theorem 5.4) for an extended statement and references. One implication is that $\pi$ generates the invariant $\sigma$-algebra on $\mathbf{X}$, so that every invariant function can be written as a measurable function of it. Items 2–3c follow directly from results of Kallenberg (2011, 2017) on the existence of universally measurable versions of $\rho$ and $P_{\tilde{\tau}|X}$. Item 3d follows from 3a and 3b.
We rely heavily on the group-orbit kernel defined in (9), where the following is claimed.
Lemma A.2.
Let $\mathbf{G}$ act properly on $\mathbf{X}$, $P_{\Gamma|[X]}$ be as defined in (9), and $P_{X}$ be any probability measure on $\mathbf{X}$. If $X\sim P_{X}$ then $(\rho(X),X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X))$.
Proof.
For any measurable $f\colon[\mathbf{X}]\times\mathbf{X}\to\mathbb{R}$, by 3(c) of Proposition A.1, standard disintegration of measure, and Fubini’s theorem,
$\int$ $f(\rho(x),x)P_{X}(dx)=\int f(\rho(x),\tilde{\tau}\rho(x))P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$=\int f([x],\tilde{\tau}[x])P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x^{\prime})P_{X\mid[X]}(dx^{\prime}\mid[x])(\rho_{*}P_{X})(d[x])$
$=\int f([x],\gamma[x])P_{\Gamma|[X]}(d\gamma\mid[x])(\rho_{*}P_{X})(d[x])$
$=\int f(\rho(x),\gamma\rho(x))P_{\Gamma|[X]}(d\gamma\mid\rho(x))P_{X}(dx)\;.$
Appendix B Proofs
B.1 Proof of Theorem 2.2
The proof of Theorem 2.2 relies on two auxiliary results. The first is a characterization of distributional invariance, which is well-studied in probability (Kallenberg, 2005). See Eaton (1989) for an accessible treatment of the basic results and some of their applications in classical statistics. We summarize two characterizations of invariance that are needed below.
Lemma B.1.
Let $\mathbf{G}$ be a compact group acting on $\mathbf{X}$ and $P$ a probability measure on $\mathbf{X}$. Let $\rho$ be a measurable orbit selector. With $X\sim P$, the following are equivalent.
  1. I0.
    $P$ is $\mathbf{G}$-invariant.
  2. I1.
    If $G\sim\lambda$ with $G{\;\perp\!\!\!\perp\;}X$, then $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GX$.
  3. I2.
    If $G\sim\lambda$ and $W\sim\rho_{*}P$ with $G{\;\perp\!\!\!\perp\;}W$, then $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GW$.
Proof.
Suppose that $P$ is $\mathbf{G}$-invariant. Then for any measurable $f\colon\mathbf{X}\to\mathbb{R}$,
$\int_{\mathbf{X}\times\mathbf{G}}f(gx)P(dx)\lambda(dg)=\int_{\mathbf{X}\times\mathbf{G}}f(x)P(dx)\lambda(dg)=\int_{\mathbf{X}}f(x)P(dx)\;,$
which shows that Property I0 implies Property I1. For the converse, assume that $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GX$. Then for any $h\in\mathbf{G}$,
$\int_{\mathbf{X}}f(hx)P(dx)$ $=\int_{\mathbf{X}\times\mathbf{G}}f(hgx)P(dx)\lambda(dg)$
$=\int_{\mathbf{X}\times\mathbf{G}}f(gx)P(dx)\lambda(dg)$
$=\int_{\mathbf{X}}f(x)P(dx)\;,$
where the second equality follows from the invariance of Haar measure.
The proof that Property I2 is equivalent to Property I0 is more involved. An accessible proof can be found in Eaton (1989, Theorems 4.3–4.4); see also Kallenberg (2017, Theorem 7.15). ∎
We note that Property I2 holds even conditionally on $\rho(X)$. That is,
$(\rho(X),X,G)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),G\rho(X),G)\;,$
which implies that ${X\mid\rho(X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)\mid\rho(X)}$.
One implication of this is that $\rho(X)$ is a sufficient statistic for the set of $\mathbf{G}$-invariant probability measures on $\mathbf{X}$ (Farrell, 1962; Dawid, 1985; Bloem-Reddy and Teh, 2020). (In fact, any maximal invariant is sufficient.) To state the result precisely, for each $x\in\mathbf{X}$, let $\Phi_{\rho(x)*}\lambda$ denote the pushforward of Haar measure by the group action applied to $\rho(x)\in\mathbf{X}$.
Corollary B.2.
Let $\mathbf{G}$ be a compact group acting on $\mathbf{X}$, and $\mathcal{P}^{\circ}(\mathbf{X})$ be the set of $\mathbf{G}$-invariant probability measures on $\mathbf{X}$. Then $\rho(X)$ is a sufficient statistic for $\mathcal{P}^{\circ}(\mathbf{X})$. That is, for each $P\in\mathcal{P}^{\circ}(\mathbf{X})$,
[…]
The second auxiliary result establishes a useful identity for exchangeable sequences of random variables, used extensively in the literature on randomization tests. A proof can be found in, e.g., Dufour (2006). For simplicity, we assume that the probability of ties is zero, but that case can be handled with a randomized tie-breaking procedure described by Dufour (2006); Hemerik and Goeman (2018).
Lemma B.3 (Dufour (2006), Proposition 2.2).
Let $S_{0},S_{1},\dotsc,S_{B}$ be an exchangeable sequence of $\mathbb{R}$-valued random variables such that $\mathbb{P}(S_{i}=S_{j})=0$ for $i\neq j$, $i,j\in\{0,\dotsc,B\}$. Define
$p_{B}=\frac{1+\sum_{b=1}^{B}\mathbbm{1}\{S_{b}\geq S_{0}\}}{B+1}\;.$
Then for any $\alpha\in[0,1]$,
$\mathbb{P}(p_{B}\leq\alpha)=\frac{\lfloor\alpha(B+1)\rfloor}{B+1}\;.$
Proof of Theorem 2.2.
Due to the sufficiency of $\rho(X)$ for $\mathcal{P}^{\circ}(\mathbf{X})$, the samples $X^{(b)}_{1:n}=(G^{(b)}_{1}X_{1},\dotsc,G^{(b)}_{n}X_{n})$ are conditionally i.i.d. (and therefore conditionally exchangeable) given $\rho(X)_{1:n}$, with the same conditional distribution as $X_{1:n}$ under $H_{0}$. Consequently, $\left(T(X^{(b)}_{1:n})\right)_{b=0}^{B}$ are conditionally i.i.d. given $\rho(X)_{1:n}$, with the same conditional distribution as $T(X_{1:n})$ under the null. It then follows from Lemma B.3 that (4) holds.
B.2 Proofs for Section 4
Proof of Theorem 4.1.
Define the random variable $\tilde{Y}\coloneqq\tilde{\tau}^{-1}Y$, where $\tilde{\tau}\sim P_{\tilde{\tau}|X}$, so that $P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}|\tilde{\tau},x)=\int\delta_{\tilde{\tau}^{-1}y}(d\tilde{y})P_{Y|X}(dy|x)$. For any integrable function $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$, by Fubini’s theorem,
$\int f(\tilde{\tau},x,\tilde{y})$ $P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$=\int f(\tilde{\tau},x,\tilde{\tau}^{-1}y)P_{Y|X}(dy\mid x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)\;.$
From this follows the identity $P_{\tilde{Y}|\tilde{\tau},X}(B|\tilde{\tau},x)=P_{Y|X}(\tilde{\tau}B|x)$.
Now assume that $P_{Y|X}$ is equivariant, so that for each $g\in\mathbf{G},x\in\mathbf{X},B\in\mathbf{S}_{\mathbf{Y}}$, $P_{Y|X}(B|gx)=P_{Y|X}(g^{-1}B|x)$ for $P_{X}$-almost all $x$. Then for any $\tilde{\tau}\in\mathbf{G},x\in\mathbf{X},g\in\mathbf{G}$ and integrable $f\colon\mathbf{Y}\to\mathbb{R}$,
$\int P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid g\tilde{\tau},gx)f(\tilde{y})$ $=\int f((g\tilde{\tau})^{-1}y)P_{Y|X}(dy\mid gx)$
$=\int f(\tilde{\tau}^{-1}g^{-1}gy)P_{Y|X}(dy\mid x)$
$=\int f(\tilde{\tau}^{-1}y)P_{Y|X}(dy\mid x)$
$=\int f(\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)\;.$
This shows that the mapping $(\tilde{\tau},x)\mapsto P_{\tilde{Y}|\tilde{\tau},X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\tilde{\tau},x)$ is $\mathbf{G}$-invariant. Therefore, by Proposition A.1, for any measurable maximal invariant $\tilde{M}\colon\mathbf{G}\times\mathbf{X}\to\mathbf{M}$, there is a unique Markov probability kernel $\tilde{R}\colon\mathbf{M}\times\mathbf{S}_{\mathbf{Y}}\to[0,1]$ such that
$P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)=P_{\tilde{Y}|\tilde{M}}(B\mid\tilde{M}(\tilde{\tau},x))\;,\quad\tilde{\tau}\in\mathbf{G},\ x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}}\;.$
Because the action of $\mathbf{G}$ on itself is transitive (i.e., there is only one orbit in $\mathbf{G}$), any maximal invariant $M$ for $\mathbf{G}$ acting on $\mathbf{X}$ is also a maximal invariant for $\mathbf{G}$ acting on $\mathbf{G}\times\mathbf{X}$, and
$P_{\tilde{Y}|\tilde{\tau},X}(B\mid\tilde{\tau},x)=\tilde{R}(B\mid M(x))\;,\quad\tilde{\tau}\in\mathbf{G},\ x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}}\;.$ (23)
This is enough to establish the desired conditional independence in (13): For any integrable $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
$\int$ $f(\tilde{\tau},x,\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$=\int f(\tilde{\tau},x,\tilde{y})\tilde{R}(d\tilde{y}\mid M(x))P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)\;.$
Conversely, assume that $(\tilde{\tau},X){\;\perp\!\!\!\perp\;}\tilde{\tau}^{-1}Y\mid M(X)$. Then (23) holds for $P_{X}\otimes P_{\tilde{\tau}|X}$-almost all $(\tilde{\tau},x)\in\mathbf{G}\times\mathbf{X}$. In particular, $P_{\tilde{Y}|\tilde{\tau},X}$ is $\mathbf{G}$-invariant in $(\tilde{\tau},x)$ for $P_{X}\otimes P_{\tilde{\tau}|X}$-almost all $(\tilde{\tau},x)$. Recall also that the inversion kernel $P_{\tilde{\tau}|X}$ is $\mathbf{G}$-equivariant. Therefore, for any integrable $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$ and any $g\in\mathbf{G}$,
$\int f(\tilde{\tau},x,y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,\tilde{\tau}(\tilde{\tau}^{-1}y))P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid g\tilde{\tau},gx)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},g^{-1}x,\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid g\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid g^{-1}x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\tilde{\tau},g^{-1}x,g^{-1}\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\tilde{\tau},g^{-1}x,g^{-1}y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\tilde{\tau},x,g^{-1}y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid gx)P_{Y|X}(dy\mid gx)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,g^{-1}y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid gx)P_{X}(dx)\;.$
This implies that
$P_{Y|X}(B\mid x)=P_{Y|X}(gB\mid gx)\;,\quad B\in\mathbf{S}_{\mathbf{Y}},\ g\in\mathbf{G},\ P_{X}\text{-a.e.\ }x\in\mathbf{X}\;.$ (24)
The subset of $\mathbf{X}$ for which (24) holds is a $\mathbf{G}$-invariant set (Kallenberg, 2017, Lemma 7.7), and therefore the possible exceptional null set on which $P_{Y|X}$ is not equivariant does not depend on $g$. If there is such an exceptional null set on which $P_{Y|X}$ is not equivariant, denoted $N^{\times}$, define the Markov kernel $Q^{\prime}$ as
$Q^{\prime}(B\mid x)\coloneqq\begin{cases}P_{Y|X}(B\mid x)&\text{if }x\notin N^{\times}\;,\\ \int_{\mathbf{G}}P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(\tilde{\tau}^{-1}B\mid\tilde{\tau}^{-1}x)&\text{if }x\in N^{\times}\;.\end{cases}$
Since $P_{\tilde{\tau}|X}$ and $P_{Y|X}$ are probability kernels, so too is $Q^{\prime}$. It is also straightforward to show that $Q^{\prime}$ is $\mathbf{G}$-equivariant, so that $Q^{\prime}$ is a regular version of $P_{Y|X}$ that is $\mathbf{G}$-equivariant for all $x\in\mathbf{X}$, and equivalent to $P_{Y|X}$ up to the null set $N^{\times}$.
If there exists a measurable representative inversion (function) $\tau$, then the same proof holds with the inversion kernel $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ substituted by $\delta_{\tau(x)}$, resulting in the simplified conditional independence statement in (12).
If the action of $\mathbf{G}$ on $\mathbf{Y}$ is trivial, then $\tilde{Y}=Y$. Moreover, $\tilde{\tau}{\;\perp\!\!\!\perp\;}Y\mid X$ by construction, and therefore $(\tilde{\tau},X){\;\perp\!\!\!\perp\;}Y\mid M(X)$ is implied by $X{\;\perp\!\!\!\perp\;}Y\mid M(X)$. ∎
For the proof of the distributional identities in Proposition 4.2 that are equivalent to equivariance, the following lemma will be useful.
Lemma B.4.
Under the conditions of Proposition 4.2, $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if $(X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\Gamma\rho(X^{\prime}),\Gamma\tilde{Y})$, where $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}X^{\prime}$ and $\Gamma,\tilde{Y}$ are sampled as in (14).
Proof.
We will prove that the distributional identity implies that $P_{Y|X}$ is equivariant. The converse follows as a special case of the stronger statements in the proof of Proposition 4.2 below.
Fix arbitrary $g\in\mathbf{G}$. From (i) the definition of $P_{\Gamma|[X]}$, (ii) Fubini’s theorem, and (iii) the equivariance of the inversion kernel $P_{\tilde{\tau}|X}$, for any integrable $f\colon\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
$\int f(x,y)P_{Y|X}(dy\mid x)P_{X}(dx)$
$=\int f(\gamma[x],\gamma\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{\Gamma|[X]}(d\gamma\mid[x])P_{[X]}(d[x])$
[…]
$=\int f(gx,gy)P_{Y|X}(dy\mid x)(g^{-1}_{*}P_{X})(dx)$
$=\int f(x,gy)P_{Y|X}(dy\mid g^{-1}x)P_{X}(dx)\;,$
which shows that $P_{Y|X}$ is equivariant up to a $P_{X}$-null set. By a similar argument as the one in the proof of Theorem 4.1, the exceptional null set is $\mathbf{G}$-invariant and therefore there is a version of $P_{Y|X}$ that is equivariant for all $g\in\mathbf{G}$, $x\in\mathbf{X}$. ∎
Proof of Proposition 4.2.
Let $(X,Y)\sim P_{X,Y}$. Assume that $P_{Y|X}$ is
$\mathbf{G}$-equivariant. To show Property E1, let $P_{[\mathbf{X}]}=\rho_{*}P_{X}$, so that for any integrable $f\colon[\mathbf{X}]\times\mathbf{Y}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
$\int$ $f(\rho(x),\tilde{y},x,y)P(dx,d\tilde{y},dy)$
$=\int f([x],\tilde{y},x^{\prime},y)P_{Y|X}(dy\mid\tilde{y},x^{\prime})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{X|[X]}(dx^{\prime}\mid[x])P_{[X]}(d[x])$
$=\int f([x],\tilde{y},\tilde{\tau}[x],\tilde{\tau}\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x^{\prime})P_{X|[X]}(dx^{\prime}\mid[x])P_{[X]}(d[x])$
$=\int f([x],\tilde{y},\gamma[x],\gamma\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{\Gamma|[X]}(d\gamma\mid[x])P_{[X]}(d[x])$
$=\int f(\rho(x),\tilde{y},\gamma\rho(x),\gamma\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid\rho(x))P_{\Gamma|[X]}(d\gamma\mid\rho(x))P_{X}(dx)\;,$
where the second equality is implied by Theorem 4.1, and the third equality follows from Fubini’s theorem. This shows that $\mathbf{G}$-equivariance of $P_{Y|X}$ implies Property E1.
To show that Property E1 implies that $P_{Y|X}$ is $\mathbf{G}$-equivariant, observe that $(\rho(X),\tilde{Y},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},\Gamma\rho(X),\Gamma\tilde{Y})$ implies that $(X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\Gamma\rho(X),\Gamma\tilde{Y})$, where $\Gamma,\tilde{Y}$ are sampled as in (14). By Lemma B.4, that implies that $P_{Y|X}$ is $\mathbf{G}$-equivariant.
Property E2 follows directly from the proof of Theorem 4.1, and Property E3 is a special case of Property E1 with $\tilde{Y}$ marginalized. ∎
Proof of Corollary 4.3.
Property E4Property E6 follow as special cases of the corresponding properties in Proposition 4.2, marginalizing over $X$. ∎
Proof of Corollary 4.4.
Properties E7 and E8 follow from the fact that
$P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|[x])=\lambda({\,\vbox{\hbox{\tiny$\bullet$}}\,})$ for each $[x]$ if and only if $P_{X}$ is $\mathbf{G}$-invariant. ∎
B.3 Proofs for Section 5
Proof of Proposition 5.1.
By Proposition 4.2, each of the randomization procedures generate conditionally i.i.d. samples, conditioned on $S_{1:n}$. By construction, the randomized samples have the same distribution as the observed samples under the null hypothesis. Then (17) follows from Lemma B.3.
The asymptotic consistency follows from the strong consistency of the empirical CDF (by the Glivenko–Cantelli theorem van der Vaart, 1998, Thm. 19.1), along with the fact that under the null hypothesis that $P_{Y|X}$ is $\mathbf{G}$-equivariant, the observations have the same distribution as the randomized samples. That is, the Glivenko–Cantelli theorem implies that for fixed $S_{1:n}$,
$\frac{1}{B+1}\sum_{b=1}^{B}\mathbbm{1}\left\{T\left((X,Y)_{1:n}^{(b)}\right)\geq T\left((X,Y)_{1:n}\right)\right\}\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{\to}}1-F_{T|S}(T(X,Y)_{1:n}\mid S_{1:n})\;,$
where $F_{T|S}$ is the conditional CDF of the test statistic given $S$ under the randomization distribution. This holds for $P$-almost every sequence $S_{1:n}$. ∎
Proof of Proposition 5.2.
For each $m$, $(X,Y)_{i,m}^{(b)}$ are independent across $i$ and conditionally independent across $b$, conditioned on $S_{1:n}$. Hence, the conditional convergence in distribution $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)_{i}$ holds jointly across $i$ and $b$. By the continuous mapping theorem, $(T((X,Y)^{(b)}_{1:n,m}))_{1:B}$ converges jointly in distribution (across $b$) to $(T((X,Y)^{(b)}_{1:n}))_{1:B}$. By assumption, ties occur with limiting probability zero and the distribution of $T((X,Y)_{1:n})$ is atomless. Therefore, again by the continuous mapping theorem, it follows that $p_{B,m}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}p_{B}$. ∎
B.4 Proofs for Section 6
B.4.1 Proof of Theorem 6.1
The proof of Theorem 6.1 builds on the analysis of Sriperumbudur et al. (2012) on the empirical estimation of MMD. In particular, for RKHS $\mathcal{H}$ and $\mathcal{H}|_{1}=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leq 1\}$, the MMD for probability measures $P,Q$ on $\mathbf{X}$ can be written as
$\mathrm{MMD}_{\sigma}(P,Q)^{2}=\left(\sup_{f\in\mathcal{H}|_{1}}\left|\int f(x)P(dx)-\int f(x)Q(dx)\right|\right)^{2}$
$=\mathbb{E}[k(X,\tilde{X})]+\mathbb{E}[k(X^{\prime},\tilde{X}^{\prime})]-2\mathbb{E}[k(X,X^{\prime})]\;,\quad X,\tilde{X}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P\;,X^{\prime},\tilde{X^{\prime}}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}Q\;,$
with empirical estimator
$\mathrm{MMD}_{\sigma}\textbf{}(\hat{P},\hat{Q})^{2}=\left(\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-f(X^{\prime}_{i}))\right|\right)^{2}\;,\quad X_{i}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P\;,\ X^{\prime}_{i}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}Q\;.$
The proof of Theorem 6.1 relies on the following lemma, which is a special case of Proposition B.1 from (Sriperumbudur et al., 2012), which itself extends a version of Talagrand’s inequality (Christmann and Steinwart, 2008, Thm. 7.5 and A.9.1) to apply to sequences of independent but not necessarily identically distributed random variables.
Lemma B.5.
Let $M\geq 0$, $n\geq 1$. For each $i\in[n]$, let $(\mathbf{X},\mathcal{A},P_{i})$ be a probability space. Let $\mathcal{F}$ be a set of $\mathcal{A}$-measurable $\mathbb{R}$-valued functions. Let $\theta_{i}\colon\mathcal{F}\times\mathbf{X}\to\mathbb{R}$ be bounded measurable functions, such that for each $f\in\mathcal{F}$, $i\in[n]$,
  • $\int_{\mathbf{X}}\theta_{i}(f,x)P_{i}(dx)=0$;
  • $\int_{\mathbf{X}}\theta_{i}(f,x)^{2}P_{i}(dx)\leq\sigma^{2}_{i}<\infty$; and
  • $\|\theta_{i}(f,{\,\vbox{\hbox{\tiny$\bullet$}}\,})\|_{\infty}\leq M<\infty$.
Define $\mathbf{W}=\mathbf{X}^{n}$, $P=P_{1}\times\dotsb\times P_{n}$, and $H\colon\mathbf{W}\to\mathbb{R}$ by
$H(w)=\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{n}\theta_{i}(f,x_{i})\right|\;,\quad w=(x_{1},\dotsc,x_{n})\;.$
Then for all $\epsilon>0$,
$P\left\{w\in\mathbf{W}:H(w)-\mathbb{E}[H(W)]\geq\epsilon\right\}\leq e^{-\frac{3}{2}\frac{\epsilon^{2}}{3a+\epsilon}}\;,$ (25)
with $a=2M\mathbb{E}[H(W)]+\sum_{i=1}^{n}\sigma^{2}_{i}$. Written differently, for all $\eta>0$,
$P\left\{w\in\mathbf{W}:H(w)\geq\mathbb{E}[H(W)]+\sqrt{2a\eta}+2\eta M/3\right\}\leq e^{-\eta}\;.$ (26)
In order to apply the lemma to our setting, observe that our MMD-based test statistic evaluated on two (conditionally) independent randomization samples $Z_{i}\sim P_{i}$ and $Z_{n+i}\sim Q_{i}$ can be written as, with $b_{i}=(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}$,
$T(\hat{P}_{1:n},\hat{Q}_{1:n})=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}b_{i}f(Z_{i})\right|=\left(\frac{1}{n^{2}}\sum_{i=1,j=1}^{2n}b_{i}b_{j}k(Z_{i},Z_{j})\right)^{1/2}\;.$ (27)
Since there is a different distribution for each $i$, consider the “population” version of the statistic to be
$T(P_{1:n},Q_{1:n})=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}b_{i}\mathbb{E}[f(Z_{i})]\right|=\left(\frac{1}{n^{2}}\sum_{i=1,j=1}^{2n}b_{i}b_{j}\mathbb{E}[k(Z_{i},Z_{j})]\right)^{1/2}\;,$ (28)
which is equal to 00 if $P_{i}=Q_{i}$ for each $i\in[n]$. Finally, observe that with
$\theta_{i}(f,x)=(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\frac{1}{n}\bigg{(}f(x)-\int f(x^{\prime})P_{i}(dx^{\prime})\bigg{)}\;,\quad f\in\mathcal{H}|_{1}\;,$ (29)
the conditions of Lemma B.5 are satisfied, with $\sigma_{i}^{2}\leq\nu^{2}/n^{2}$ and $M\leq 4\nu/n$.
We will use it to bound
$|T(\hat{P}_{1:n},\hat{Q}_{1:n})-T(P_{1:n},Q_{1:n})|$
$\qquad\leq\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}(f(Z_{i})-\mathbb{E}[f(Z_{i})])\right|$
$\qquad=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}\theta_{i}(f,Z_{i})\right|\;.$
We will make repeated use of the following upper bound on the expectation of $H(W)$, a proof of which can be found in, for example, (Sriperumbudur et al., 2012, Appendix D).
Lemma B.6.
Let $\theta_{i}$ be as defined in (29) and $\mathcal{F}=\mathcal{H}|_{1}$. Then
$\mathbb{E}[H(W)]\leq 2\nu\sqrt{\frac{2}{n}}\;.$
The factor of $2$ comes from a symmetrization argument (Sriperumbudur et al., 2012, Appendix B.2). The rest of the upper bound is a bound on the Rademacher complexity of the RKHS.
Proof of Theorem 6.1.
Let $P_{i}$ be the conditional distribution of $Z_{i}$ conditioned on $S_{i}=s_{i}$, and let $Q_{i}=P_{i}$ be the same for $Z^{\prime}_{i}$. In particular, with $\hat{\Delta}_{n}=|T(\hat{P},\hat{Q}^{(0)})|$ and $\bar{H}_{n}=\mathbb{E}[H(W)]$, if $\hat{\Delta}_{n}>\bar{H}_{n}$ then
$p(\hat{\Delta}_{n})\coloneq\;$ :- $\mathbb{P}\left(\left|T(\hat{P}^{(b)},\hat{Q}^{(b)})\right|\geq\hat{\Delta}_{n}\mid S_{1:n},\hat{\Delta}_{n}\right)$
$\leq\exp\left(-\frac{3}{2}\frac{(\hat{\Delta}_{n}-\bar{H}_{n})^{2}}{3a+\hat{\Delta}_{n}-\bar{H}_{n}}\right)$
$\leq\exp\left(-\frac{3}{2}(\hat{\Delta}_{n}-\bar{H}_{n})+\frac{9}{2}a\right)$
$\leq\exp\left(-\frac{3}{2}\hat{\Delta}_{n}+\frac{9}{2n}\nu^{2}+\frac{16\nu^{2}\sqrt{2}}{\sqrt{n^{3}}}+3\nu\sqrt{\frac{2}{n}}\right)=:\hat{p}(\hat{\Delta}_{n})\;,$
where the first inequality follows from Lemma B.5, the second inequality follows from $e^{-1/(1+x)}\leq e^{-(1-x)}$, $x\geq 0$, and the third inequality follows from $\sigma^{2}_{i}\leq\nu^{2}/n^{2}$ and Lemma B.6.
Now, conditioned on $\hat{\Delta}_{n}$ and $S_{1:n}$, $N_{B}\coloneq p_{B}(B+1)-1$ is a random variable with $\textrm{Binomial}(B,p(\hat{\Delta}_{n}))$ distribution, so that
$\mathbb{P}(p_{B}\leq\alpha$ $\mid S_{1:n},\hat{\Delta}_{n})=\mathbb{P}(N_{B}\leq\alpha(B+1)-1\mid S_{1:n},\hat{\Delta}_{n})$
$=F_{B,p(\hat{\Delta}_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
where $F_{B,p(\hat{\Delta}_{n})}$ is the $\textrm{Binomial}(B,p(\hat{\Delta}_{n}))$ distribution function. It is known that for $q\geq p$, a $\textrm{Binomial}(B,q)$-distributed random variable $X_{B,q}$ stochastically dominates a $\textrm{Binomial}(B,p)$-distributed random variable $X_{B,p}$, in the sense that
$\mathbb{P}(X_{B,q}\geq k)\geq\mathbb{P}(X_{B,p}\geq k)\;,\quad\text{all }k\in[B]\;.$
This implies the reverse ordering for the CDFs, so that the bound $p(\hat{\Delta}_{n})\leq\hat{p}(\hat{\Delta}_{n})$ implies
$\mathbb{P}(p_{B}\leq\alpha$ $\mid S_{1:n},\hat{\Delta}_{n})\geq F_{B,\hat{p}(\hat{\Delta}_{n})}(\lfloor\alpha(B+1)-1\rfloor)$
$=\sum_{k=0}^{\lfloor\alpha(B+1)-1\rfloor}\binom{B}{k}\hat{p}(\hat{\Delta}_{n})^{k}(1-\hat{p}(\hat{\Delta}_{n}))^{B-k}\;.$
Now, applying Lemma B.5 to (27) with $\hat{P}_{1:n},Q_{1:n}$ the empirical distributions of the observations $Z_{1:n}$ and the randomizations $Z_{1:n}^{(0)}$, respectively, conditioned on $S_{1:n}$, we have that with probability at least $1-e^{-\eta}$,
$\left|T(\hat{P}_{1:n},\hat{P}^{(0)}_{n})-T(P_{1:n},P^{(0)}_{1:n})\right|\leq 2\nu\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{2\frac{\nu^{2}}{n}+32\nu^{2}\sqrt{\frac{2}{n^{3}}}}+\frac{8\eta\nu}{3n}\;.$
Therefore, letting $\Delta_{n}\coloneq T(P_{1:n},P^{(0)}_{1:n})$ (as in (28)), with probability at least $1-e^{-\eta}$,
$\hat{\Delta}_{n}\geq\Delta_{n}-2\nu\sqrt{\frac{2}{n}}-\sqrt{\eta}\sqrt{2\frac{\nu^{2}}{n}+32\nu^{2}\sqrt{\frac{2}{n^{3}}}}-\frac{8\eta\nu}{3n}\;,$
Putting this all together, with probability at least $1-e^{-\eta}$,
$\mathbb{P}(p_{B}\leq\alpha$ $\mid S_{1:n})\geq\sum_{k=0}^{\lfloor\alpha(B+1)-1\rfloor}\binom{B}{k}p_{\eta}(\Delta_{n})^{k}(1-p_{\eta}(\Delta_{n}))^{B-k}\;,$
where
$p_{\eta}(\Delta_{n})=\exp\left(-\frac{3}{2}\Delta_{n}+R_{n}(\eta)\right)\;,$
with
$R_{n}(\eta)\coloneq\frac{9}{2n}\nu^{2}+\frac{16\nu^{2}\sqrt{2}}{\sqrt{n^{3}}}+6\nu\sqrt{\frac{2}{n}}+\frac{3\nu}{2}\sqrt{\eta}\sqrt{\frac{2}{n}+32\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta\nu}{n}\;.$
B.4.2 Proof of Corollary 6.2
Proof of Corollary 6.2.
Let $B_{\eta}(\Delta_{n})=Bp_{\eta}(\Delta_{n})$. By the Chernoff bound for binomial random variables,
$\mathbb{P}(N_{B}\leq B_{\alpha}\mid S_{1:n})\geq 1-\exp\left(-B_{\alpha}\ln\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}+B_{\alpha}-B_{\eta}(\Delta_{n})\right)\;.$
This is at least $\beta$ if and only if
$-B_{\alpha}\ln\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}+B_{\alpha}-B_{\eta}(\Delta_{n})\leq\ln(1-\beta)\;.$ (30)
Now,
$\ln\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}=-\ln\left(1-\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)=\sum_{k=1}^{\infty}\frac{1}{k}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{k}\;,$
and therefore (30) becomes
$B_{\alpha}\sum_{k=1}^{\infty}\frac{1}{k}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{k}-(B_{\alpha}-B_{\eta}(\Delta_{n}))$
$=B_{\alpha}\sum_{k=2}^{\infty}\frac{1}{k}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{k}\geq-\ln(1-\beta)\;.$
The inequality is satisfied if
$B_{\alpha}\frac{1}{2}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{2}\geq-\ln(1-\beta)\;.$
Solving for $\Delta_{n}$ yields
$\Delta_{n}\geq\frac{2}{3}\left(R_{n}(\eta)-\ln\frac{B_{\alpha}}{B}-\ln\left(1-\sqrt{\frac{-2\ln(1-\beta)}{B_{\alpha}}}\right)\right)\;.$
B.5 Power lower bounds for adaptive statistics
As with Theorem 6.1, the conditional probability of rejecting when using the FUSE statistic can be lower bounded in terms of
$\bar{\Delta}_{n}\coloneq\frac{1}{L}\sum_{\ell=1}^{L}\sqrt{\mathbb{E}\left[\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z^{\prime(0)}_{1:n})\mid S_{1:n}\right]}\;.$
For convenience, let $R_{n}^{\vee}(\eta)$ denote the maximum over $\ell\in[L]$ of $R_{n,\ell}(\eta)$, which is obtained by substituting $\nu_{\ell}$ for $\nu$ into (19).
Proposition B.7.
Let $T$ be the FUSE statistic using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. Assume that for each $\ell\in\{1,\dotsc,L\}$, $\Delta_{n,\ell}\geq 2\nu_{\ell}\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-Le^{-\eta}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\bar{\Delta}_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
where $F_{B,p_{\eta}(\Delta_{n}^{\vee})}$ is the distribution function for the binomial distribution of $B$ trials with success probability
$p_{\eta}(\bar{\Delta}_{n})=\min\left\{1,L\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+R_{n}^{\vee}(\eta)\right)\right\}\;.$
Using this, a lower bound for $\bar{\Delta}_{n}$ that guarantees power $\beta$ with high probability, analogous to the one in Corollary 6.2, can be obtained.
For the following proof, define $\nu_{\vee}\coloneq\sup_{\ell}\nu_{\ell}$, so that
$a^{\vee}\coloneq\sup_{\ell\in[L]}a_{\ell}\leq\frac{\nu_{\vee}^{2}}{n}\left(1+16\sqrt{\frac{2}{n}}\right)\;,$
and similarly $\bar{H}^{\vee}\coloneq\sup_{\ell}\bar{H}_{\ell}\leq 2\nu_{\vee}\sqrt{2/n}$. Finally, let
$R_{n}^{\vee}(\eta)\coloneq\frac{9}{2n}\nu_{\vee}^{2}+\frac{16\nu_{\vee}^{2}\sqrt{2}}{\sqrt{n^{3}}}+6\nu_{\vee}\sqrt{\frac{2}{n}}+\frac{3\nu_{\vee}}{2}\sqrt{\eta}\sqrt{\frac{2}{n}+32\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta\nu_{\vee}}{n}\;.$
Proof of Proposition B.7.
Using the V-statistic to estimate MMD, recall that
$\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},Z_{1:n}^{\prime})=\frac{1}{\omega}\log\left(\frac{1}{L}\sum_{\ell=1}^{L}\exp\left(\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)\right)\;.$
Therefore, using the fact that if $\mathbb{E}[X_{1}]\geq\mathbb{E}[X_{2}]$ then $\sup X_{1}\geq\mathbb{E}[X_{2}]$, denoting $\mathcal{A}_{1:n}=(S_{1:n},Z_{1:n},Z^{(0)}_{1:n})$,
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\mathbb{P}\left(\sup_{\ell}\exp\left[\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\right]\geq\right.$
$\qquad\qquad\left.\frac{1}{L}\sum_{\ell=1}^{L}\exp\left[\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},{Z}_{1:n}^{(0)})\right]\mid\mathcal{A}_{1:n}\right)$
$=\mathbb{P}\left(\sup_{\ell}\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)\;.$
By a Fréchet inequality and Lemma B.5,
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\bigg{|}\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\bigg{|}\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}\mid\mathcal{A}_{1:n}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\frac{\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})}{3a_{\ell}+\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}}\right)\right\}$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}+\frac{9}{2}a^{\vee}+\frac{3}{2}\bar{H}^{\vee}\right)\right\}\;.$
Now by Jensen’s inequality,
$\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},Z_{1:n}^{\prime})=\frac{1}{\omega}\log\left(\frac{1}{L}\sum_{\ell=1}^{L}\exp\left(\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)\right)$
$\geq\frac{1}{\omega}\frac{1}{L}\sum_{\ell=1}^{L}\log\exp\left(\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)$
$\geq\left(\frac{1}{L}\sum_{\ell=1}^{L}\bigg{|}\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\bigg{|}\right)^{2}\;.$
Therefore,
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\frac{1}{L}\sum_{\ell=1}^{L}\bigg{|}\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},{Z}_{1:n}^{(0)})\bigg{|}+\frac{9}{2}a^{\vee}+\frac{3}{2}\bar{H}^{\vee}\right)\right\}\;.$
Thus, for any $\eta>0$, with probability at least $1-Le^{-\eta}$,
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+R^{\vee}_{n}(\eta)\right)\right\}\;.$
The rest of the proof is analogous to the proof of Theorem 6.1. ∎
When using the supremum kernel as the test statistic, we expect that when
$\Delta_{n}^{\vee}=\sup_{\ell\in[L]}\Delta_{n,\ell}\coloneq\sup_{\ell\in[L]}\sqrt{\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\mid S_{1:n}\right]}$
is large, then the test should reject with high probability.
Proposition B.8.
Let $T$ be the supremum kernel distance using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. Assume that for each $\ell\in\{1,\dotsc,L\}$, $\Delta_{n,\ell}\geq 2\nu_{\ell}\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-Le^{-\eta}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\Delta_{n}^{\vee})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
where $F_{B,p_{\eta}(\Delta_{n}^{\vee})}$ is the distribution function for the binomial distribution of $B$ trials with success probability
$p_{\eta}(\Delta_{n}^{\vee})=\min\left\{1,L\exp\left(-\frac{3}{2}\Delta_{n}^{\vee}+R_{n}^{\vee}(\eta)\right)\right\}\;.$
Using this, a lower bound for $\Delta_{n}^{\vee}$ that guarantees power $\beta$ with high probability, analogous to the one in Corollary 6.2, can be obtained.
Proof of Proposition B.8.
Recall that using the V-statistic to estimate the MMD,
$\widehat{\mathrm{SK}}(Z_{1:n},Z_{1:n}^{\prime})$ $=\sup_{\ell\in\{1,\dotsc,L\}}\left\{\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right\}$
$=\sup_{\ell\in\{1,\dotsc,L\}}\left\{\left|\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right|\right\}^{2}\;.$
Therefore, by a Fréchet inequality and Lemma B.5, and denoting
$\mathcal{A}_{1:n}=(S_{1:n},Z_{1:n},Z^{(0)}_{1:n})$,
$\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$=1-\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\leq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\left|\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\right|\leq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}\mid\mathcal{A}_{1:n}\right)\right\}$
$=\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\frac{\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})}{3a_{\ell}+\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}+\frac{9}{2}a_{\ell}+\frac{3}{2}\bar{H}_{\ell}\right)\right\}\;.$
Using $a^{\vee}$ and $\bar{H}^{\vee}$ as defined above,
$\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid S_{1:n},Z_{1:n},{Z}_{1:n}^{(0)}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}+\frac{9}{2}a^{\vee}+\frac{3}{2}\bar{H}^{\vee}\right)\right\}\;.$
Thus, for any $\eta>0$, with probability at least $1-Le^{-\eta}$,
$\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid S_{1:n}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\Delta_{n}^{\vee}+R_{n}^{\vee}(\eta)\right)\right\}\;,$
where $\Delta_{n}^{\vee}=\sup_{\ell\in[L]}\Delta_{n,\ell}$, and $R_{n}(\eta)$ is as defined above. The rest of the proof is analogous to the proof of Theorem 6.1. ∎
Appendix C Examples of exact conditional randomization
Here, we describe two possible assumptions for $P_{\Gamma|[X]}$ that each allow for sampling from the exact conditional distribution.
Assumption 1.
$P_{\Gamma|[X]}=P_{\Gamma}$ for all $M(X)$.
Assumption 1 says that the distribution over the group action is the same across orbits. Generating a new sample $\mathsf{\Gamma}_{1},\dotsc,\mathsf{\Gamma}_{n}$ can then be done by simply permuting $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$. The resulting randomization procedure using Property E3 is outlined in Algorithm 4. We note that this assumption is trivially true if $\mathbf{G}$ acts transitively on $\mathbf{X}$.
Algorithm 4 $\mathsf{\Gamma}$ randomization procedure under Assumption 1
1:sample $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$ with $\tilde{\tau}_{i}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X_{i})$
2:sample permutation $\tilde{\tau}_{\pi(1)},\dotsc,\tilde{\tau}_{\pi(n)}$
3:compute $\left((\tilde{\tau}_{\pi(1)}\rho(X_{1}),\tilde{\tau}_{\pi(1)}\tilde{Y}_{1}),\dotsc,(\tilde{\tau}_{\pi(n)}\rho(X_{n}),\tilde{\tau}_{\pi(n)}\tilde{Y}_{n})\right)$
Assumption 2.
Suppose there is a group $\mathbf{H}$ acting transitively on $\mathbf{M}$, and let $\phi\colon\mathbf{H}\to\mathbf{G}$ be a group homomorphism. Then $P_{\Gamma|[X]}$ is $\mathbf{H}/\mathbf{G}$-equivariant in the sense that for all $h\in\mathbf{H}$,
$P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid hM(X))=P_{\Gamma|[X]}(\phi(h)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid M(X))\;.$
Assumption 2 says that the conditional distribution over the group action is itself equivariant with respect to the action of $\mathbf{H}$ on $\mathbf{M}$. Assuming that $\phi$ is known, new samples can be generated by permuting $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$ and applying a corresponding transform under $\phi$. The corresponding randomization procedure using Property E3 is outlined in Algorithm 5.
Algorithm 5 $\mathsf{\Gamma}$ randomization procedure under Assumption 2
1:compute $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$ with $\tilde{\tau}_{i}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X_{i})$
2:sample permutation $\tilde{\tau}_{\pi(1)},\dotsc,\tilde{\tau}_{\pi(n)}$
3:sample $\tilde{h}_{\pi(1)},\dotsc,\tilde{h}_{\pi(n)}$ where $\phi(\tilde{h}_{\pi(i)})^{-1}\rho(X_{\pi(i)})\stackrel{{\scriptstyle\text{\rm\tiny a.s.}}}{{=}}\rho(X_{i})$
4:compute $\left((\phi(\tilde{h}_{\pi(i)})^{-1}\tilde{\tau}_{\pi(i)}\rho(X_{i}),\phi(\tilde{h}_{\pi(i)})^{-1}\tilde{\tau}_{\pi(i)}\tilde{Y}_{i})\right)_{i=1}^{n}$
Appendix D Additional experiments and experimental details
D.1 Test for invariance p-values
Figure 9 shows histograms of the $p$-values for the tests for invariance described in Example 1 from Section 2.2. In particular, the distributions of $p$-values for the randomization tests are approximately uniform.
Figure 9: Histograms showing the distribution of $1000$ $p$-values for the tests described in Example 1 when $\theta$ is sampled uniformly from $[0,p)$. Labels are the same as in Figure 2.
D.2 Additional results for the experiment in Section 7.2
For the first experiment discussed in Section 7.2, Figure 10 shows the estimated rejection rates for the tests that use independent comparison sets. For the second experiment where non-$\mathrm{SO}(d)$-equivariance is isolated to the first two dimensions, Figure 11 shows the estimated rejection rates for tests that use independent comparison sets. In both experiments, results for the tests that use independent comparison sets are similar to that of tests that reuse a set.
Figure 10: Rejection rates for tests for $\mathrm{SO}(d)$-equivariance that use independent comparison sets in the changing covariance experiment for varying values of the $Y|X$ covariance matrix off-diagonal values $p$ and sample size $n$. Dotted lines show the corresponding rates for tests that reuse a comparison set.
Figure 11: Rejection rates for tests that use independent comparison sets in the experiment where the non-SO(d)-equivariance comes from the covariance $p$ between the first two dimensions, for varying values of the sample size $n$. Dotted lines show the corresponding rates for tests that reuse a comparison set.
D.3 Effect of the number of randomizations
In the same setting described in Section 7.2 that takes the conditional distribution to be non-equivariant in all dimensions, we investigate the effect of the number of randomizations $B$ used to estimate the $p$-value. Figure 12 shows the estimated rejection rates for the equivariant and a non-equivariant setting as $B$ ranges between $50$ and $1000$. In the equivariant setting, the rejection rates remain well-under the expected rate of $0.05$. In the non-equivariant setting, the power remains relatively consistent after $B=100$ or so.
Figure 12: Rejection rates for the changing covariance experiment for number of randomizations $B$ ranging from $50$ to $1000$. For each set of tests, the conditional distribution is $\mathrm{SO}(3)$-equivariant for the plots in the first row and not equivariant for the plots in the second.
D.4 Sampling from approximation versus true conditional distribution
Consider a setting where $X$ is generated in the following way. Let $Z\sim\chi_{1}$ be a chi random variable with $1$ degree of freedom, and take $X\coloneq Z\Gamma\mathbf{e}_{1}$, where $\Gamma|Z\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|Z1_{d}+\varepsilon)$ is a $d$-dimensional rotation matrix, with $1_{d}$ being the vector of ones of length $d$ and $\varepsilon\sim\mathsf{N}(0,\mathbf{I}_{d})$. In other words, the conditional distribution $P_{\Gamma|[X]}$ is known and can be sampled from exactly. Let $Y|X$ have the conditional distribution described in Section 7.2, with $p$ being the covariance between any pair of dimensions of $Y$. Recall that the conditional distribution of $Y|X$ is $\mathrm{SO}(d)$-equivariant if and only if $p=0$.
For $d\in\{3,4,5\}$, we compare the performance of tests that sample from an approximation of $P_{\Gamma|[X]}$ as per Algorithm 2 with tests that sample from $P_{\Gamma|[X]}$ exactly. The test results are shown in Figure 13. In this setting, we see that kernel methods well-approximate sampling from the true conditional. Histograms showing the distribution of $p$-values for these tests are provided in Figures 14 and 15. In the equivariant setting $(p=0)$, we see that the distributions of $p$-values for the conditional randomization tests are approximately uniform for both tests that use approximate sampling and tests that use exact sampling.
Figure 13: Rejection rates for $\mathrm{SO}(d)$-equivariance tests that reuse a comparison set (top) and that use independent comparison sets (bottom) for varying values of the $Y|X$ covariance matrix off-diagonal values $p$ and sample size $n$. The dotted lines represent tests that sample from the true conditional $P_{\Gamma|[X]}$, and the solid lines represent tests that sample from an approximation.
Figure 14: Histograms showing the distribution of $1000$ $p$-values of the $\mathrm{SO}(3)$-equivariant tests that reuse a comparison set for the approximate versus exact sampling experiment at sample size $n=250$. Parameter $p$ is the covariance between any two dimensions.
Figure 15: Histograms showing the distribution of $1000$ $p$-values of the $\mathrm{SO}(3)$-equivariant tests that use independent comparison sets for the approximate versus exact sampling experiment at sample size $n=250$. Parameter $p$ is the covariance between any two dimensions.
D.5 Sensitivity to non-equivariance in the mean
For $d=3$, we consider two other forms of non-equivariance under the action of $\mathrm{SO}(d)$ that could manifest in the conditional distribution. Consider the conditional distribution
$Y\mid X\sim\begin{cases}\mathsf{N}(X+s1_{d},\mathbf{I}_{d})&\text{if }\cos^{-1}\left(\frac{X_{1}}{\sqrt{X_{1}^{2}+X_{2}^{2}}}\right)<p\\ \mathsf{N}(X,\mathbf{I}_{d})&\text{otherwise}\end{cases}\;.$
Non-equivariance under rotations about the origin is controlled by two parameters with this conditional distribution:
  1. 1.
    A proportion $p$ of rotations that leads to a shift in the mean, determined by the 2D angle between the vector spanned by the first two entries of $X$ and the axis of the first dimension. The conditional distribution is equivariant at both $p=0$ and $p=1$, and most non-equivariant at $p=0.5$.
  2. 2.
    The scaling factor $s$ in the shift in the mean, with the conditional distribution being equivariant only when $s=0$.
For various combinations of $p$ and $s$ and sample sizes ranging from $25$ to $500$, the test rejection rates are shown in Figure 16. Our tests are sensitive to both parameters, with powers increasing as the conditional distribution moves away from equivariance in either parameter. When a comparison set is reused, the randomization tests outperform the baseline tests in terms of power.
Figure 16: Rejection rates for tests that reuse a comparison set (top) and that use independent comparison sets (bottom) in the non-equivariant mean experiment, controlled by proportion $p$ and shift $s$ parameters, for varying values of sample size $n$. The conditional distribution is not $\mathrm{SO}(3)$-equivariant in all settings shown.
D.6 Permutation equivariance
We consider a synthetic example where we test for equivariance with respect to the action of a different group. Consider the same $X$ marginal distribution as in Section 7.2, but now consider the conditional distribution
$Y\mid X\sim\begin{cases}\mathsf{N}(X+s\mathbf{e}_{1},\mathbf{I}_{d})&\text{if }X_{1}>X_{j}\;,\;\forall j\in\{2,\dotsc,d\}\\ \mathsf{N}(X,\mathbf{I}_{d})&\text{otherwise}\end{cases}\;,$
where $s$ is some positive scaling factor. In words, the conditional distribution has a shift in the mean if the first dimension of $X$ is its largest entry. For data generated under this setup, we test for equivariance with respect to the symmetric group $\mathbb{S}_{d}$—the group of $d$-permutations. The conditional distribution is $\mathbb{S}_{d}$-equivariant if and only if $s=0$.
In this setting, a convenient maximal invariant to condition on is the order statistic $M(X)=(X_{(1)},\dotsc,X_{(d)})$. For $d\in\{3,4,5\}$ and various values of $s$, the test rejection rates are shown in Figure 17. The results are similar to that of the other synthetic experiments, though we note that the powers of the tests rapidly decreases with increasing dimensions in this particular setting.
Figure 17: Rejection rates for tests that reuse a comparison set (top) and that use independent comparison sets (bottom) for the permutation equivariance experiment for varying values of sample size $n$. The conditional distribution is $\mathbb{S}_{d}$-equivariant only at $s=0$.
D.7 Distribution of transverse momenta in LHC dijet data
Figure 18 shows the distribution of transverse momenta for a sample of $500$ first and second leading constituents from the LHC dijet data. Both distributions visually appear to be $\mathrm{SO}(2)$-invariant.
Figure 18: Transverse momentum of $500$ first and second leading constituents (LC) in the LHC dijet data.
{acks}
[Acknowledgments] The authors would like to thank Arthur Gretton and Antonin Schrab for a helpful conversation about randomization tests and MMD estimators.
{funding}
This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of British Columbia. BBR and KC are supported by the Natural Sciences and Engineering Research Council of Canada (NSERC): RGPIN2020-04995, RGPAS-2020-00095, DGECR-2020-00343. AS is supported by a Canadian Statistical Sciences Institute (CANSSI) Distinguished Postdoctoral Fellowship.